### v1

# Support Recovery and $\ell_{2}$-Error Bound for Sparse Regression with Quadratic Measurements via Weakly-Convex-Concave Regularization

Abstract
The recovery of unknown signals from quadratic measurements finds extensive applications in fields such as phase retrieval, power system state estimation, and unlabeled distance geometry. This paper investigates the finite sample properties of weakly convex–concave regularized estimators in high-dimensional quadratic measurements models. By employing a weakly convex–concave penalized least squares approach, we establish support recovery and $\ell_{2}$-error bounds for the local minimizer. To solve the corresponding optimization problem, we adopt two proximal gradient strategies, where the proximal step is computed either in closed form or via a weighted $\ell_{1}$ approximation, depending on the regularization function. Numerical examples demonstrate the efficacy of the proposed method.
Keywords: Nonconvex statistics, Finite sample error bound, Consistency, Optimization algorithm.
1 Introduction
We consider the following quadratic measurement regression model
$y_{i}=\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}+\varepsilon_{i},\quad i=1,\ldots,n,$ (1)
where $y_{i}\in\mathbb{R}$ denotes the observed response, $\bm{Z}_{i}\in\mathbb{R}^{d\times d}$ is a symmetric design matrix, $\bm{\beta}\in\mathbb{R}^{d}$ is the vector of unknown parameters representing true signals, and $\varepsilon_{i}\in\mathbb{R}$ is the random noise.
Model (1) arises in numerous applications in physics, engineering, and data science, including phase retrieval (Candes et al. 2015), generalized phase retrieval (Wang and Xu 2019), the unassigned distance geometry problem (Huang and Dokmanić 2021), and power system state estimation (Wang et al. 2019). Recent years have witnessed a surge of interests and methodological developments for this model. Notable advances include Thaker et al. (2020), Huang et al. (2020), Chen and Ng (2022), and Fan et al. (2025a), who proposed various algorithms and theoretical guarantees of optimal solutions under various problem settings. It is worth noting that when the design matrix can be decomposed as a outer product of a real vector $\bm{Z}_{i}=\bm{z}_{i}\bm{z}_{i}^{\top}$, model (1) is referred to as real-valued phase retrieval. In this case, it also represents a special case of the single index model, with index $\bm{z}_{i}^{\top}\bm{\beta}$ and square link function. Single index models are widely used in statistics and econometrics (Xu et al. 2022, Yang et al. 2024).
In high-dimensional regime where $d$ is large or even exceeds $n$, a common assumption is that the true signal is sparse, i.e., only a small number of its entries are non-zero. This sparsity assumption is crucial for ensuring identifiability and statistical efficiency, leading to the problem of sparse regression with quadratic measurements. Several studies have explored this challenging problem. For instance, Fan et al. (2018) studied an $\ell_{q}(0<q<1)$-regularized least squares method, established a weak oracle property and proposed a fixed-point iterative algorithm. Bolte et al. (2018) called the regularized least squares problem for model (1) quadratic inverse problem and proposed a Bregman proximal gradient algorithm (PGA). Subsequently, Zhang et al. (2023) and Ding et al. (2025) developed Bregman PGA to solve regularized problems including quadratic inverse problem as a special case. While existing literature has primarily focused on algorithmic development for the regularized least squares formulation of quadratic inverse problem (or model (1)), the corresponding statistical guarantees remain relatively underexplored. More recently, Chen et al. (2025) employed the thresholded Wirtinger flow algorithm originally proposed by (Cai et al. 2016) for sparse phase retrieval to solve the regularized least squares problem for model (1). A special case of sparse quadratic model is sparse phase retrieval that attracted considerable attention (Cai et al. 2016, 2022, Soltanolkotabi 2019, Xia and Xu 2021, Huang and Xu 2024).
To address sparse regression in linear models, various other regularization techniques have been developed, such as the $\ell_{q}(0<q<1)$ penalty (Frank and Friedman 1993), the SCAD (Fan and Li 2001), and the MCP (Zhang 2010). It is well-known that LASSO has constant derivative that induces persistent shrinkage and biased estimation(Zou 2006). In contrast, concave penalties (e.g., SCAD, MCP, and $\ell_{q}(0<q<1)$) have derivatives that vanish for large coefficients, thereby mitigating shrinkage bias and yielding nearly unbiased estimates. However, nonconvex penalties introduces computational challenges. In particular, highly nonconvex penalties may cause numerous local minima, and the convergence of first-order algorithm (e.g., PGA) relies on favorable properties of the objective function. One such property is weak convexity, which is a key condition that generally ensures the proximal operator is single-valued (Wang 2010, Khanh et al. 2025), under standard technical assumptions. If this condition is violated, the proximal operator may be multi-valued, which poses a potential risk that iterative algorithms could fail to converge to a meaningful solution.
This challenge motivates the study of regularizers that balance statistical and computational considerations. Loh and Wainwright (2015) developed a general statistical analysis for a class of weakly convex regularized M-estimators and established finite sample error bounds between any stationary point of the penalized estimator and the population-optimal solution without regularization (i.e., the solution minimizing the expected population risk). While their framework applies to a wide range of models, including linear and generalized linear models and graphical LASSO, it cannot be directly applied to sparse quadratic measurement problems. The main difficulty is that the least squares loss for this model leads to a highly nonconvex quartic polynomial that are computationally very challenging to minimize. In fact, as pointed out by Candes et al. (2015), it is already known that determining whether a stationary point of a quartic polynomial is a local minimizer is NP-hard. Moreover, Loh and Wainwright (2015) imposed an additional constraint, which introduces an extra tuning parameter that must be chosen carefully to ensure that the true value $\bm{\beta}^{*}$ is a feasible point, thereby further increasing the difficulty of solving the problem. Weakly convex regularization methods have also been widely used in sparse signal recovery (Yang et al. 2019, Komuro et al. 2022) and linear inverse problems (Shumaylov et al. 2024, Goujon et al. 2024, Ebner et al. 2025), but these works do not explicitly exploit concave penalties.
To mitigate the above mentioned numerical issues, in this paper, we propose a specific class of concave penalties that also satisfy weak convexity, termed Weakly-Convex-Concave Penalties (WCCP). The major advantage of this class is that it provides a flexible framework for analyzing local minimizers in sparse quadratic measurement problems, combining the statistical benefits of concavity with the algorithmic guarantees of weak convexity.
Specifically, in this paper, we study the regularization problem
$\min_{\bm{\beta}\in\mathbb{R}^{d}}~\frac{1}{4n}\sum_{i=1}^{n}\left(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i}\right)^{2}+P_{\lambda_{n}}(\bm{\beta}),$ (2)
where $\lambda_{n}>0$ is a tuning parameter that encourages sparsity, and $P_{\lambda_{n}}(\cdot)$ is a WCCP. As mentioned before, model (1) constitutes a special case of the single index model with a known square link function. Therefore, the proposed method (2) has potential applications for single index models.
Our contributions are threefold. First, we bridge a notable gap in the literature by providing the first systematic statistical analysis of the WCCP-regularized estimator for the quadratic measurement model (1). We establish rigorous support recovery guarantees and $\ell_{2}$-error bounds for its local minimizer. we derive the statistical properties of the WCCP-regularized estimator for the quadratic measurement model (1) and establish support recovery and tight $\ell_{2}$-error bounds for the local minimizer. Second, we develop iterative algorithms tailored to different regularization functions that provide convergence guarantees. Third, we demonstrate the effectiveness of our method, particularly in the context of sparse phase retrieval, and show that it performs comparable or better than the existing LASSO-based approaches.
The rest of this paper is organized as follows. Section 2 presents the assumptions and statistical properties of the estimator. Section 3 introduces two iterative algorithms and provides a convergence analysis. Numerical experiments are presented in Section 4. Technical lemmas and proofs are deferred to Appendices A and B.
Notation
For a vector $\bm{v}=(v_{1},\ldots,v_{d})^{\top}$, let $\|\bm{v}\|$, $\|\bm{v}\|_{1}$, and $\|\bm{v}\|_{\infty}$ denote its Euclidean, $\ell_{1}$, and $\ell_{\infty}$ norms, respectively. Denote by $\bm{e}_{d,j}$ the $j$th column of the $d\times{}d$ identity matrix $\bm{I}_{d}$. For any $n\times{}d$ matrix $\bm{Z}$, denote $|\bm{Z}|_{\infty}=\max_{1\leq i\leq n,1\leq j\leq{}d}|\bm{e}_{n,i}^{\top}\bm{Z}\bm{e}_{d,j}|$. The submatrix of $\bm{Z}$ with rows and columns indexed by sets $\Gamma_{1}$ and $\Gamma_{2}$ is denoted $\bm{Z}^{\Gamma_{1},\Gamma_{2}}$. The subvector of $\bm{\beta}$ indexed by $\Gamma$ is denoted $\bm{\beta}_{\Gamma}$. The $j$th standard basis vector in $\mathbb{R}^{d}$ is denoted $\bm{e}_{d,j}$. For a function $h(t)$, we denote its gradient and Hessian by $\nabla h(t)$ and $\nabla^{2}h(t)$, respectively, and its subdifferential by $\partial h(t)$. Let $\bm{\beta}^{\star}$ be the true parameter value, and denote $\Gamma^{\star}=\mathrm{supp}(\bm{\beta}^{\star}):=\{j:{\bm{e}_{d,j}^{T}}\bm{\beta}^{\star}\neq 0,j=1,...,d\}$. Define the oracle regularized least square estimator as follows,
$\hat{\bm{\beta}}_{1}^{o}\in\arg\min_{\bm{\beta}_{1}\in\mathbb{R}^{s}}\frac{1}{4n}\sum\limits_{i=1}^{n}({{\bm{\beta}_{1}}^{\top}\bm{Z}_{i}^{\Gamma^{\star}\Gamma^{\star}}{\bm{\beta}_{1}}-y_{i}})^{2}+\sum_{j\in\Gamma^{\star}}p_{\lambda_{n}}(|\beta_{j}|).$
2 Finite Sample Statistical Results
In this section, we derive the finite sample properties and consistency of the weakly convex-concave estimator in the high-dimensional case where $d$ is large than $n$. We assume that $\ln d=o(n^{\alpha})$ with some constant $\alpha\in(0,1)$ and denote the number of non-zero elements of the true signal by $s$. Following literature (Huang et al. 2008, Fan et al. 2018), we assume that there exists constants $0<c_{1}\leq c_{2}<\infty$ such that
$c_{1}\leq\min\{|\bm{e}_{d,j}^{\top}\bm{\beta}^{\star}|,j\in\Gamma^{\star}\}\leq\max\{|\bm{e}_{d,j}^{\top}\bm{\beta}^{\star}|,j\in\Gamma^{\star}\}\leq c_{2}.$ (3)
Further regarding the observed response and design matrix, following Fan et al. (2018), the data are assumed to be standardized as
$\sum\limits_{i=1}^{n}y_{i}=0,~~~\mbox{and}~~~\sum\limits_{i=1}^{n}|\bm{Z}_{i}|_{\infty}^{2}=n.$ (4)
For the regularization function, noise and design matrix, we make the following assumptions.
Assumption 1.
The regularization function is coordinate separable:
$P_{\lambda_{n}}(\bm{\beta})=\sum_{j=1}^{d}p_{\lambda_{n}}(|\bm{\beta}_{j}|),$
for some scalar function $p_{\lambda_{n}}$ which satisfies
  • (i)
    The function $p_{\lambda_{n}}$ is concave on $(0,\infty)$ and there exists $\mu>0$ such that $p_{\lambda_{n}}(t)+\frac{\mu}{2}t^{2}$ is convex.
  • (ii)
    The function $p_{\lambda_{n}}$ satisfies $p_{\lambda_{n}}(0)=0$ and $p_{\lambda_{n}}(t)>0$ for $t>0$.
  • (iii)
    For $t>0$, the function $p_{\lambda_{n}}$ is non-decreasing and $t\mapsto\frac{p_{\lambda_{n}}(t)}{t}$ is non-increasing in $t$.
  • (iv)
    The function $p_{\lambda_{n}}$ is differentiable with derivative $p^{\prime}_{\lambda_{n}}$ for all $t>0$, and subdifferentiable at $t=0$, satisfying $\lim_{t\to 0^{+}}p^{\prime}_{\lambda_{n}}(t)=\lambda_{n}\varrho$ with $\varrho>0$. The function $p^{\prime}_{\lambda_{n}}$ is locally Lipschitz continuous in $(0,\infty)$.
Compared to the conditions of (Loh and Wainwright 2017, Loh 2017), Assumption 1 additionally requires the penalty function to be concave and its derivative $p_{\lambda_{n}}^{{}^{\prime}}$ to be locally Lipschitz continuous on the positive half-line. In fact, the condition used by (Loh and Wainwright 2017, Loh 2017) is relaxed enough to include the LASSO penalty. As mentioned earlier, concave penalties are known to possess stronger variable selection capabilities. Therefore, we explicitly emphasize the concavity of the penalty function in our framework. Compared to a general concave penalty, the imposition of weak convexity here constrains its degree of nonconvexity. As previously mentioned, this ensures the proximal mapping of the penalty function remains single-valued, which is crucial for achieving superior algorithmic convergence. Combining the second order continuous differentiability of the loss and the assumption about the locally Lipschitz’s continuity of $p_{\lambda_{n}}^{{}^{\prime}}$ enables us to use the generalized Hessian to verify that a stationary point is a local minimizer. In fact, many concave penalties, including SCAD and MCP, possess locally Lipschitz continuous first derivatives on the positive half-line.
Moreover, in addition to the well-known SCAD and MCP regularizers, several other commonly used regularizers also satisfy the conditions in Assumption 1, we present a few examples below.
Example 1. Firm regularizer (Woodworth and Chartrand 2016):
$p_{\lambda_{n}}(t)=\left\{\begin{array}[]{rcl}\lambda_{n}|t|-\frac{\lambda_{n}t^{2}}{2c},&{|t|<c},\\ \frac{c\lambda_{n}}{2},&{|t|\geq c},\end{array}\right$.
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=1$ and $\mu=\frac{1}{c}$.
Example 2. LOG regularizer (Lobo et al. 2007):
$p_{\lambda_{n}}(t)=\lambda_{n}\log(c|t|+1),$
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=\frac{1}{c}$ and $\mu=c$.
Example 3. EXP regularizer (Bradley et al. 1998):
$p_{\lambda_{n}}(t)=\lambda_{n}(1-e^{-c|t|}),$
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=c$ and $\mu=c^{2}$.
Assumption 2.
Let errors $\{\varepsilon_{1},\cdots,\varepsilon_{n}\}$ be independent and identically distributed Sub-Gaussian with variance proxy $\sigma^{2}$, and $\varepsilon_{i}$ have zero mean and positive variance $E\varepsilon_{i}^{2}$.
Assumption 3.
For any $\bm{x},\bm{y}\in\mathbb{R}^{s}$, there exist constants $c_{4}\geq c_{3}>0$ satisfying that
$c_{3}{\|\bm{x}\|}^{2}{\|\bm{y}\|}^{2}\leq\frac{1}{n}\sum\limits_{i=1}^{n}(\bm{x}^{\top}\bm{Z}^{\Gamma^{\star}\Gamma^{\star}}_{i}\bm{y})^{2}\leqslant c_{4}{\|\bm{x}\|}^{2}{\|\bm{y}\|}^{2}.$
To establish the finite‑sample properties of our proposed estimator, we first analyze the oracle estimator $\hat{\bm{\beta}}_{1}^{o}$, which is defined with prior knowledge of the true support set. The following theorem provides its non‑asymptotic consistency guarantee.
Theorem 1.
Under model (1) and Assumption 1-3, if $c_{4}\geq 3/(ns)$, then
$\mathbb{P}\left(\|\hat{\bm{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\leq r_{n}\right)\geq 1-p_{1}-p_{2},$
where $p_{1}=2\exp\left\{-s\ln(1+2n)\right\},p_{2}=\frac{1}{n^{-n/2}}$,
$r_{n}=C_{0}\left(\sqrt{\ln(1+2n)/n}+\lambda_{n}\varrho/\sqrt{s}\right)$ and $C_{0}=\max\left\{10\sigma\sqrt{c_{4}}/c_{1}c_{3},4/c_{1}^{2}c_{3}\right\}$.
The condition $c_{4}\geq 3/(ns)$ is non‑restrictive, as the right‑hand side diminishes to near zero for any practical $n$ and $s$, meaning any finite uniform bound $c_{4}$ will automatically satisfy it. Theorem 1 shows that the estimation error of the oracle estimator is bounded by $r_{n}$. Under the standard high‑dimensional scaling $(s\ln n+\ln d)/n\to 0$, one can choose $\lambda_{n}$ such that $\lambda_{n}/\sqrt{s}\to 0$, which ensures $r_{n}\to 0$ and thus the consistency of $\hat{\bm{\beta}}_{1}^{o}$ as $n\to\infty$.
Building on this oracle result, a standard and necessary step for analyzing the high-dimensional estimator is to impose conditions on the design matrix $\bm{Z}_{i}$ to control its column correlations.
Assumption 4.
There exists a constant $c_{5}>0$ such that
$|\sum\limits_{i=1}^{n}\bm{Z}_{i}^{\Gamma^{\star}\Gamma^{\star}}\otimes\bm{Z}_{i}^{\Gamma^{\star c}\Gamma^{\star c}}|_{\infty}\leq c_{5}\sqrt{n},~|\sum\limits_{i=1}^{n}\bm{Z}_{i}^{\Gamma^{\star}\Gamma^{\star}}\otimes\bm{Z}_{i}^{\Gamma^{\star c}\Gamma^{\star}}|_{\infty}\leq c_{5}\sqrt{n},$
where $\otimes$ is the Kronecker product.
The inequalities in Assumption 4 are similar to the partial orthogonality condition proposed by Huang et al. (2008) for linear models, which was later extended to quadratic models by Fan et al. (2018). They can be regarded as that the sub-matrix of $\bm{Z}_{i}$ corresponding to $\Gamma^{\star}$ and the other complementary matrix being orthogonal.
Theorem 2.
Under model (1) and Assumption 1-4, suppose
$36c_{2}^{2}c_{5}C_{0}s^{2}\leq\sqrt{n},\quad\sqrt{\frac{2\ln(1+2n)}{n}}\leq c_{2}C_{0}^{-1},\quad\frac{3}{\sqrt{ns}}\leq c_{2}$ (5)
and
$\max\Big\{\sqrt{\frac{s\ln(1+2n)}{n}},6\sqrt{2}c_{2}\sigma s\sqrt{\frac{s\ln(1+2n)+\ln d}{n}}\Big\}\leq\lambda_{n}\varrho\leq\frac{c_{2}\sqrt{s}}{\sqrt{2}C_{0}}.$ (6)
If $\mu\leq c_{1}^{2}c_{3}s$ and
$2(4c_{3}+3c_{4})c_{2}r_{n}+10\sigma\sqrt{c_{4}}\sqrt{\frac{\ln(1+2n)}{n}}\leq c_{1}^{2}c_{3}\sqrt{s},$ (7)
then there exists a local minimizer $\hat{\bm{\beta}}$ of (2) such that
  • (i)
    $\mathbb{P}(\hat{\bm{\beta}}_{{\Gamma^{\star}}^{c}}=0)\geq 1-p_{1}-2p_{2}-\frac{2}{n}$,
  • (ii)
    $\mathbb{P}\left(\min\{\|\hat{\bm{\beta}}+\bm{\beta}^{\star}\|,\|\hat{\bm{\beta}}-\bm{\beta}^{\star}\|\}\leq r_{n}\right)\geq 1-p_{1}-2p_{2}-\frac{2}{n}$.
Remark 1.
The conditions in Theorem 2 define a non‑asymptotic feasible region for the problem dimensions $(n,s,d)$ and the regularization parameter $\lambda_{n}$, within which the desired statistical guarantees hold with explicit constants. From an asymptotic perspective ($n\to\infty$, allowing $s,d\to\infty$), these conditions are naturally satisfied under standard sparsity assumptions.
The inequalities in (5) impose constraints on the scaling between $n$ and $s$. The first inequality, $36c_{2}^{2}c_{5}C_{0}s^{2}\leq\sqrt{n}$, implies $s=O(n^{1/4})$. The remaining two inequalities involve terms that decay to zero as $n$ grows and are thus satisfied for sufficiently large $n$.
The admissible range for $\lambda_{n}$ in (6) requires that the lower bound (which dominates the noise) does not exceed the upper bound (which preserves the signal). Under the standard high‑dimensional scaling $(s\ln n+\ln d)/n\to 0$ and the sparsity condition $s=O(n^{1/4})$, the lower bound is $o(1)$ while the upper bound is $O(n^{1/8})$. Hence, for large $n$, the admissible interval for $\lambda_{n}$ is non‑empty. Moreover, by choosing $\lambda_{n}$ to be of the same order as the lower bound (which tends to zero), we ensure $\lambda_{n}\to 0$ and $\lambda_{n}/\sqrt{s}\to 0$. The upper bound is only needed to control the bias in the non‑asymptotic analysis and does not conflict with $\lambda_{n}\to 0$ asymptotically.
The condition $\mu\leq c_{1}^{2}c_{3}s$ restricts the weak convexity parameter $\mu$ of the penalty function, ensuring that it does not dominate the curvature of the loss function. For typical weakly convex penalties (e.g., SCAD, MCP), $\mu$ is a fixed constant. If $s\to\infty$, the right‑hand side grows unbounded, so a fixed $\mu$ satisfies the inequality for large $n$; if $s$ remains bounded, the condition imposes a mild fixed upper bound on $\mu$.
The inequality (7) involves $r_{n}=C_{0}\bigl(\sqrt{\ln(1+2n)/n}+\lambda_{n}\varrho/\sqrt{s}\bigr)$. Because $\lambda_{n}/\sqrt{s}\to 0$ as argued above, we have $r_{n}\to 0$ as $n\to\infty$, which implies the estimation error bound in Theorem 2 vanishes asymptotically.
In summary, these conditions jointly describe a regime where sparsity grows slowly and dimension may grow exponentially with $n$, a typical setting in high‑dimensional statistics. Therefore, all conditions in Theorem 2 are compatible with both finite‑sample guarantees and asymptotic consistency.
Remark 2.
For the sparse phase retrieval problem, Cai et al. (2016) considered parameter estimation of real-valued signals by minimizing the empirical $\ell_{2}$ loss function. Their results show that the estimator $\hat{\bm{\beta}}$ satisfies $\min\{\|\hat{\bm{\beta}}+\bm{\beta}^{\star}\|,\|\hat{\bm{\beta}}-\bm{\beta}^{\star}\|\}=O(\sqrt{\ln d/n})$ as $\ln d/n\rightarrow 0$. In our result, when the regularization parameter satisfies $\lambda_{n}=O(\sqrt{(s\ln n+\ln d)/n})$, the error bound simplifies to $O(\sqrt{\ln d/n})$ under the assumption that $s\ln(n)=O(\ln d)$, which shows that our estimator attains the known optimal rate for sparse phase retrieval. Recently, Huang and Xu (2020) studied the estimation performance of the nonlinear Lasso of phase retrieval. Their results show that the estimator $\hat{\bm{\beta}}$ satisfies $\min\{\|\hat{\bm{\beta}}+\bm{\beta}^{\star}\|,\|\hat{\bm{\beta}}-\bm{\beta}^{\star}\|\}\leq\|\varepsilon\|/\sqrt{n}$. According to Cai et al. (2009), the error $\bm{\varepsilon}$ satisfies $\|\bm{\varepsilon}\|\leq\sigma\sqrt{n+2\sqrt{n\ln n}}$ with high probability as $n\rightarrow\infty$. In this case, the error bound simplifies to $\sigma\sqrt{1+2\sqrt{n\ln n}/n}$. It is clear that the estimation error exceeds the estimation error bound established in our work.
3 Optimization Algorithm
In this section, we discuss the numerical computation of the problem to assess the performance of the proposed method. Since $n$ and $d$ are given, to simplify notation we denote $\lambda_{n}=\lambda$, so that (2) is written as
$\min_{\beta\in\mathbb{R}^{p}}F(\bm{\beta}):=L(\bm{\beta})+P_{\lambda}(\bm{\beta}),$ (8)
where $L(\bm{\beta}):=\frac{1}{4n}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}$ and $\lambda>0$. Noting the term $L(\bm{\beta})$ is smooth and the term $P_{\lambda}(\bm{\beta})$ is weakly convex-concave, we employ the PGA to solve the problem. The PGA is an established classical algorithm (Beck 2017) and has been applied in various inverse and optimization problems (e.g., Bolte et al. (2018), Soltanolkotabi (2019),Zhang et al. (2023), Fan et al. (2025b)). The core component of PGA is the proximal operator. Given a function $P:\mathbb{R}^{d}\to(-\infty,\infty)$, recall that the proximal operator associated with $P$ is defined as
$\mathrm{prox}_{P}(\bm{v})=\arg\min_{\bm{u}}\left\{\frac{1}{2}\|\bm{u}-\bm{v}\|^{2}+P(\bm{u})\right\}.$
In particular, given a weight vector $\bm{w}=(w_{1},\dots,w_{d})^{\mathsf{T}}$ with $w_{j}>0$ for all $j$, we denote the proximal operator of the weighted $\ell_{1}$-norm $\|\bm{w}\circ\bm{u}\|_{1}$ at $\bm{v}$ by
$\mathcal{S}(\bm{v},\bm{w}):=\bm{v}-\max\{-\bm{w},\min\{\bm{v},\bm{w}\}\},$
where the $\max$ and $\min$ operations are performed component-wise. Note that this reduces to the classical soft thresholding operator $\mathrm{prox}_{\lambda\|\cdot\|_{1}}(\bm{v})$ when $w_{j}=\lambda$ for all $j$. For proximal operators that does not admit a closed-form solution, we can adopt weighted $\ell_{1}$ algorithm based on the concavity of the regularization function.
This analysis yields two alternative fixed-point characterizations of the minimizers as follows.
Proposition 1.
There exists a constant $r>0$, for any $\tau\in(0,\hat{L}^{-1}]$, with $\hat{L}=\sup_{\|\bm{\beta}\|\in B_{r}}\|\nabla^{2}L(\bm{\beta})\|$ and $B_{r}:=\left\{\bm{\beta}\in\mathbb{R}^{d}:\|\bm{\beta}\|\leq r\right\}$, such that for any minimizer $\hat{\bm{\beta}}$ of problem (2), it satisfies
[…]
and
$\hat{\bm{\beta}}=\mathcal{S}(\hat{\bm{\beta}}-\tau\nabla L(\hat{\bm{\beta}}),\tau\hat{\bm{w}}_{\lambda}),$ (10)
where the $j$th element of the vector $\hat{\bm{w}}_{\lambda}$ is defined as
$\bm{e}_{d,j}^{\top}\hat{\bm{w}}_{\lambda}=\begin{cases}\max\big(p_{\lambda}^{{}^{\prime}}(|\hat{\beta}_{j}|),\epsilon_{1}\big),&\text{if }\hat{\beta}_{j}\neq 0,\\ \lambda\varrho,&\text{if }\hat{\beta}_{j}=0.\end{cases}$
Here $\epsilon_{1}$ is any positive number.
These two characterizations provide the foundation for our algorithmic design. According to the relationship (9), the iterative scheme for solving (2) is outlined in Algorithm 1. For (10), the weight $\hat{\bm{w}}_{\lambda}$ depends on the unknown optimal solution $\hat{\bm{\beta}}$ rendering the problem circular and not directly solvable. To overcome this, we adopt the classic framework of iteratively reweighted $\ell_{1}$ algorithms (Zou (2006),Bai et al. 2024). Within this framework, we employ the Majorization-Minimization (MM) technique. At each iteration, the regularization term is linearly approximated via its first-order expansion, which constructs a surrogate function. This leads to a sequence of subproblems of the form
$\min_{\bm{\beta}\in\mathbb{R}^{d}}L(\bm{\beta})+\|\bm{w}_{\lambda}^{k}\circ\bm{\beta}\|_{1},$
where $\bm{w}_{\lambda}^{k}$ is adaptively updated weights computed from the current iterate $\bm{\beta}^{k}$. The next iterate $\bm{\beta}^{k+1}$ is then updated according to the rule specified in Algorithm 2, which stems from an approximate minimization step for this surrogate function.
Algorithm 1
1:Data $\{y_{i},\bm{Z}_{i}\}$, parameters $T,\epsilon>0,\delta>0,\gamma_{0}\in(0,1),\gamma_{1}\in(0,1)$
[…]
3:Initialization: Choose a spectral initialization point $\bm{\beta}^{0}$ and set $k=0$
4:Computation: Compute $\bm{\beta}^{k+1}=\text{prox}_{\tau_{k}P_{\lambda}}(\bm{\beta}^{k}-\tau_{k}\nabla L(\bm{\beta}^{k}))$, where $\tau_{k}=\gamma_{1}\gamma_{0}^{j_{k}}$ and $j_{k}$ is the smallest nonnegative integer such that:
$F(\bm{\beta}^{k})-F(\bm{\beta}^{k+1})\geq\delta\|\bm{\beta}^{k+1}-\bm{\beta}^{k}\|^{2}.$ (11)
5:while $\|\bm{\beta}^{k+1}-\bm{\beta}^{k}\|\geq\epsilon\max\{1,\|\bm{\beta}^{k}\|\}$ and $k\leq T$ do
6: Set $k=k+1$, go back to Step 3
7:end while
8:Output: $\bm{\beta}^{k+1}$
Algorithm 2
1:Data $\{y_{i},\bm{Z}_{i}\}$, parameters $T,\epsilon,\epsilon_{1}>0,\delta>0,\gamma_{0}\in(0,1),\gamma_{1}\in(0,1)$
[…]
3:Initialization: Choose a spectral initialization point $\bm{\beta}^{0}$ and set $k=0$
4:Weight computation: Take $\bm{w}_{\lambda}^{k}=(w_{1}^{k},\cdots,w_{d}^{k})^{\top}$ with $w_{\lambda,j}^{k}=\max(p_{\lambda}^{{}^{\prime}}(|\beta_{j}^{k}|),\epsilon_{1})$ if $\beta_{j}^{k}\neq 0$ and $w_{j}^{k}=\lambda\varrho$ if $\beta_{j}^{k}=0$
5:Update: Compute $\bm{\beta}^{k+1}=S(\bm{\beta}^{k}-\tau_{k}\nabla L(\bm{\beta}^{k}),\lambda\tau_{k}\bm{w}^{k})$, where $\tau_{k}=\gamma\alpha^{j_{k}}$ and $j_{k}$ is the smallest nonnegative integer such that condition (11) holds
6:while $\|\bm{\beta}^{k+1}-\bm{\beta}^{k}\|\geq\varepsilon\max\{1,\|\bm{\beta}^{k}\|\}$ and $k\leq T$ do
7: Set $k=k+1$, go back to Step 3
8:end while
9:Output: $\bm{\beta}^{k+1}$
It is worth noting that computing the step size $\tau_{k}$ is a crucial step, and its selection depends on the parameter $\hat{L}$, which is difficult to determine. To address this, we employ the Armijo line search method (11). The existence of such a smallest $j_{k}$ and the admissible range for $\tau_{k}$ are established in Lemma 5 of Appendix B.
Having established the step size selection strategy, we now turn to the convergence analysis of the algorithms.
Proposition 2.
Assume that $\left\{\bm{\beta}^{k}\right\}$ is the sequence generated by Algorithm 1 or 2. Then the following conclusions hold,
  • (i)
    $\left\{F(\bm{\beta}^{k})\right\}$ and $\left\{F^{k}(\bm{\beta}^{k})\right\}$ both are monotonically decreasing.
  • (ii)
    $\lim\limits_{k\rightarrow\infty}\|\bm{\beta}^{k+1}-\bm{\beta}^{k}\|=0.$
  • (iii)
    Every accumulation point of the sequence $\{\bm{\beta}^{k}\}$ generated by Algorithm 1 satisfies the fixed-point equation (9), and that generated by Algorithm 2 satisfies (10).
4 Numerical Experiments
In this section, we examine the finite sample performance of the proposed method through numerical simulations. All experiments are implemented in MATLAB(R2023b) on a laptop with 32GB memory. The MATLAB code is available at https://github.com/fjmath/sparseQMR_WCCP.
We compares the performance of the SCAD, MCP, Firm, LOG, and EXP regularizers with the $\ell_{1}$ and $\ell_{1/2}$ regularizers. The initial point is obtained using the sparse spectral initialization method proposed by Chen et al. (2025). For all regularization methods, the optimal step size is determined using the Armijo line search. The regularization parameter is chosen according to the scheme of Chen and Ng (2022), defined as $\lambda_{n}=\sqrt{\frac{c\ln(d)}{n^{2}}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}}\|\bm{\beta}\|$, where the coefficient $c$ is determined via cross-validation. The relative error in the experiments is computed as follows,
$\text{Relerr}=\frac{\min\left\{\|\hat{\bm{\beta}}-\bm{\beta}^{\star}\|,\|\hat{\bm{\beta}}+\bm{\beta}^{\star}\|\right\}}{\|\bm{\beta}^{\star}\|}.$
Experiment 1. The sample size is $n$, and the measurement matrix $\bm{Z}$ is taken to be a symmetric matrix with Gaussian entries. The noise $\bm{\varepsilon}$ follows a sub-Gaussian distribution. The true signal $\bm{\beta}^{\star}$ is generated as a Gaussian random vector. This experiment compares the performance of WCCP-regularized with the $\ell_{1}$ and $\ell_{1/2}$ regularizers under varying $n/d$ ratios. In the experiments, noise $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$ is added to the measurements $y_{i}$. Regarding the signal dimension $n$ and sparsity level $s$, we consider two cases: d=128,s=10d=128,~s=10 and d=256,s=15d=256,~s=15. For $\sigma=0.01,~\text{and}~0.05$, the recovery of the unknown signal $\bm{\beta}^{\star}$ is considered successful when the error is below $10^{-3}~\text{and}~10^{-2}$ respectively. The measurement ratios are set as $n/d=\left\{0.1,\dots,0.9,1\right\}$, with each experiment run 100 times. The performance is evaluated based on success rate and average relative error, with the latter is calculated only for successful trials. Furthermore, Supplementary Tables S1–S8 confirm that as $n/d$ approaches 1, several WCCP regularizers—particularly SCAD and MCP—achieve lower relative errors and higher stability than $\ell_{1}$ and $\ell_{1/2}$, while being orders of magnitude faster. In contrast, Firm, LOG and EXP incur substantially higher computational costs. Overall, SCAD and MCP offer the best trade‑off between accuracy and efficiency.
Figure 1 illustrates the curves of success rate as functions of the ratio $n/d$. The results confirm that WCCP methods generally outperform $\ell_{1}$ regularization and are competitive with $\ell_{1/2}$. The SCAD and MCP penalties achieve the highest success rates, especially as $n/d$ increases. Other WCCP members (Firm, LOG, and EXP) also exhibit strong performance, with LOG and EXP proving particularly robust across noise levels. Detailed mean squared error and variance results for all measurement ratios are provided in Appendix C (Tables 5-12). These results show that as $n/d$ approaches 1, our works exhibits superior relative-error performance. Therefore, our method is more stable in identifying the sparse support.
Figure 1:
Experiment 2. This experiment demonstrates the performance of WCCP-regularized method in the sparse phase retrieval problem. The proximal operator of the EXP regularizer is computationally expensive; therefore, Algorithm 2 is used to solve the problem in this experiment. The sparse test signal is generated from the standard test image "cameraman.tif" (available in MATLAB’s image processing toolbox) following the wavelet-based sparsification procedure Jagatap and Hegde (2018) and Shah and Hegde (2021). Our implementation adapts the publicly available code from Shah and Hegde (2021): we resize the image to 64×64 pixels, apply a 4-level Haar wavelet transform, and retain only the top $5\%$ of the coefficients by magnitude. This differs from Shah and Hegde (2021) in both the test image used and the sparsity criterion (percentage-based vs. fixed count). In the experiments, noise $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$ is added to the measurements.
The comprehensive numerical experiments demonstrate the clear advantages of WCCP-regularized methods over traditional approaches for sparse phase retrieval. As shown in Tables 1 and 3, WCCP methods, particularly SCAD and MCP, consistently achieve near-perfect reconstruction (SSIM $\approx$ 0.9992) across all noise levels and measurement ratios, while traditional $\ell_{1}$ and $\ell_{0.5}$ methods show significant performance degradation, especially under high noise conditions and low measurement ratios. Notably, MCP achieves perfect reconstruction at measurement ratio $n/d=0.4$, whereas $\ell_{1}$ requires $n/d\geq 0.9$ to reach comparable quality, demonstrating the superior sparse recovery capability of WCCP regularizers.
In addition to enhanced reconstruction quality, WCCP methods offer substantial computational advantages. As evidenced in Tables 2 and 4, SCAD and MCP complete reconstruction tasks in significantly less time compared to $\ell_{1}$ regularization, while maintaining superior accuracy. For instance, at noise level $\sigma=10^{-4}$, $\ell_{1}$ requires 139.08 seconds while SCAD and MCP complete in only 0.61 and 0.58 seconds respectively. These results confirm that WCCP-regularized methods provide an efficient and robust framework for sparse signal recovery, making them particularly suitable for practical applications requiring high-quality reconstruction under challenging conditions.
Figure 2:
Table 1: Structural Similarity Index (SSIM) performance of sparse regularization methods at different noise levels (measurement ratio $n/d=0.50$)
Method Noise Level $\sigma$
0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ 1
L1 0.8142 (0.3295) 0.8410 (0.3246) 0.8020 (0.3508) 0.8805 (0.2901) 0.8416 (0.3238) 0.8825 (0.2848) 0.7733 (0.3554)
SCAD 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
MCP 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
$L_{0.5}$ 0.9597 (0.1765) 0.9582 (0.1831) 0.9614 (0.1690) 0.9992 (0.0000) 0.9197 (0.2446) 0.9577 (0.1855) 0.9221 (0.2373)
Firm 0.8787 (0.2960) 0.9564 (0.1912) 0.7615 (0.3731) 0.8071 (0.3422) 0.8418 (0.3231) 0.9206 (0.2424) 0.8782 (0.2956)
LOG 0.9212 (0.2401) 0.9992 (0.0000) 0.9992 (0.0000) 0.9191 (0.2466) 0.9591 (0.1792) 0.9992 (0.0000) 0.9591 (0.1793)
EXP 0.8353 (0.3372) 0.8310 (0.3453) 0.7186 (0.3931) 0.8468 (0.3129) 0.8371 (0.3337) 0.9122 (0.2675) 0.7116 (0.4023)
Table 2: Computational time of sparse regularization methods at different noise levels (measurement ratio $n/d=0.50$)
Method Noise Level $\sigma$
0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ 1
L1 1.7725 1.9901 139.0842 1.9897 1.8820 1.9666 1.3740
SCAD 0.5458 0.6096 0.6105 0.5703 0.5774 0.5958 0.7209
MCP 0.5206 0.5516 0.5760 0.5341 0.5313 0.7673 0.6821
$L_{0.5}$ 0.4745 0.5401 0.5185 0.5150 0.4807 0.4364 0.3899
Firm 0.6950 0.7791 0.7060 0.6844 0.6630 84.0347 0.5769
LOG 1.1945 1.3752 1.3729 1.1638 1.2105 1.3171 0.9131
EXP 1.1392 15.2622 1.1028 1.1917 1.1054 1.2542 0.8009
Table 3: Structural Similarity Index (SSIM) performance at different measurement ratios (noise level $\sigma=10^{-2}$)
Method Measurement Ratio $(n/d)$
0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
L1 0.1255 (0.0708) 0.4185 (0.2080) 0.6918 (0.3482) 0.8556 (0.2946) 0.8875 (0.2736) 0.8916 (0.2631) 0.9647 (0.1541) 0.9992 (0.0000) 0.9646 (0.1543)
SCAD 0.4886 (0.0312) 0.6165 (0.0409) 0.9082 (0.1958) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
MCP 0.4614 (0.0298) 0.5978 (0.0347) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
$L_{0.5}$ 0.2317 (0.1600) 0.6062 (0.0509) 0.9019 (0.2536) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
Firm 0.0820 (0.0823) 0.5535 (0.3544) 0.7897 (0.3724) 0.9112 (0.2708) 0.9287 (0.2172) 0.9648 (0.1535) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
LOG 0.2770 (0.1481) 0.5740 (0.1302) 0.8487 (0.2430) 0.9180 (0.2500) 0.9614 (0.1689) 0.9259 (0.2256) 0.9992 (0.0000) 0.9992 (0.0000) 0.9992 (0.0000)
EXP 0.0564 (0.0198) 0.1998 (0.1355) 0.7809 (0.3571) 0.7959 (0.3620) 0.8439 (0.3187) 0.8873 (0.2733) 0.9292 (0.2156) 0.9992 (0.0000) 0.9992 (0.0000)
Table 4: Computational time at different measurement ratios (noise level $\sigma=10^{-2}$)
Method Measurement Ratio $(n/d)$
0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
L1 0.8037 1.5078 1.9668 1.9321 2.0649 2.0567 2.1940 2.3576 2.3357
SCAD 0.3875 0.4539 0.5693 0.5984 0.6111 0.6245 0.6540 0.6864 0.7105
MCP 0.3673 0.4219 0.5328 0.5630 0.5906 0.5945 0.6095 0.6428 0.6673
$L_{0.5}$ 0.3269 0.3634 0.4327 0.4728 0.5130 0.5303 0.5673 0.5848 0.6059
Firm 0.4306 0.5050 0.6558 0.6647 0.6916 0.7159 0.7635 0.7762 0.8061
LOG 0.7325 0.8526 1.0285 1.1950 1.3259 1.4074 1.5055 1.5480 1.6134
EXP 0.7063 0.8373 1.0086 1.1550 1.2877 1.3680 1.4670 1.5155 1.5763
[…]
Without loss of generality, let $\Gamma^{\star}=\{1,...,s\}$ and $\bm{\beta}^{\star}=(\bm{\beta}_{1}^{\star\top},\bm{0}^{\top})^{\top}$, and then denote
$\bm{Z}_{i}=\begin{bmatrix}\bm{Z}_{i}^{11}&\bm{Z}_{i}^{12}\\ \bm{Z}_{i}^{21}&\bm{Z}_{i}^{22}\end{bmatrix}$.
For convenience, we denote
$F(\bm{\beta}):=L(\bm{\beta})+P_{\lambda_{n}}(\bm{\beta}),$
where $L_{n}(\bm{\beta}):=\frac{1}{4n}\sum\limits_{i=1}^{n}({{\bm{\beta}}^{\top}\bm{Z}_{i}{\bm{\beta}}-y_{i}})^{2}$. Since $p_{\lambda}^{\prime}(t)$ is locally Lipschitz continuous on $(0,\infty)$, by Rademacher’s theorem, the second derivative $p_{\lambda}^{\prime\prime}(t)$ exists almost everywhere. For any $t_{0}>0$, the Clarke generalized second-order derivative of $p_{\lambda}$ at $t_{0}$ is defined as
$\partial^{2}p_{\lambda}(t_{0})=\operatorname{co}\Big\{\lim_{k\to\infty}p_{\lambda}^{\prime\prime}(t_{k}):t_{k}\to t_{0},\;p_{\lambda}^{\prime\prime}(t_{k})\ \text{exists}\Big\},$
where $\operatorname{co}$ denotes the convex hull (all convex combinations of the set). Moreover, since $p_{\lambda}$ is $\mu$-weakly convex, its Clarke generalized second-order derivative satisfies
$\partial^{2}p_{\lambda}(t)\subset[-\mu,+\infty),\quad\forall\,t>0.$
The following lemma is analogous to Lemma 3.1 in Fan et al. (2025a) but uses a different tail bound for the noise term, which leads to a slightly different probability lower bound and a different constant in the deviation inequality.
Lemma 1.
Under Assumptions 2-4, if $n\geq 3$ and $c_{4}\geq 3/(ns)$ it follows that
$\mathbb{P}(E^{1})\geq 1-2\exp\{-d\ln(1+2n)\}-\frac{1}{n^{-n/2}},$
where the event is defined by
$E^{1}:=\{\sup_{\|\bm{u}\|=1,\|\bm{v}\|=1}|\frac{1}{n}\sum_{i=1}^{n}\bm{u}^{\top}\bm{Z}_{i}^{11}\bm{v}\varepsilon_{i}|\leq 5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}\}.$
Proof.
The proof follows the same structure as Lemma 3.1 in Fan et al. (2025a). The only difference is our choice of the parameter $t$ when bounding $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}$ via Bernstein’s inequality. While Fan et al. (2025a) uses $t=\sigma^{2}$, we take $t=16\sigma^{2}\ln n$. This yields a different tail probability $n^{-n/2}$ and, under the condition $c_{4}\geq 3/(ns)$, leads to the deviation inequality
$\sup_{\|\bm{u}\|=1,\|\bm{v}\|=1}\Bigl|\frac{1}{n}\sum_{i=1}^{n}\bm{u}^{\top}\bm{Z}_{i}^{11}\bm{v}\varepsilon_{i}\Bigr|\leq 5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}.$
We now detail the modified steps. By Assumption 2 and Lemma 1.12 of Rigollet and Hütter (2017), $\varepsilon_{i}^{2}-\mathbb{E}\varepsilon_{i}^{2}$ is sub‑exponential with parameter $16\sigma^{2}$. Applying Bernstein’s inequality (Theorem 1.13 in Rigollet and Hütter (2017)) with $t=16\sigma^{2}\ln n$, we obtain
$\mathbb{P}\Bigl(\frac{1}{n}\sum_{i=1}^{n}\bigl(\varepsilon_{i}^{2}-\mathbb{E}\varepsilon_{i}^{2}\bigr)>16\sigma^{2}\ln n\Bigr)\leq\exp\bigl\{-n\ln n/2\bigr\}=n^{-n/2}.$
Since $\sigma^{2}\geq\mathbb{E}\varepsilon_{1}^{2}$ (due to the sub‑Gaussian property), it follows that
$\mathbb{P}\Bigl(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}>\sigma^{2}+16\sigma^{2}\ln n\Bigr)\leq n^{-n/2}.$ (12)
The remaining arguments are identical to those in Fan et al. (2025a). In particular, with probability at least $1-2\exp\{-d\ln(1+2n)\}-n^{-n/2}$,
$\sup_{\|\bm{u}\|=1,\|\bm{v}\|=1}\Bigl|\frac{1}{n}\sum_{i=1}^{n}\bm{u}^{\top}\bm{Z}_{i}^{11}\bm{v}\varepsilon_{i}\Bigr|\leq\sigma\sqrt{6c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}+\frac{\sqrt{\sigma^{2}+16\sigma^{2}\ln n}}{n}.$
Using $c_{4}\geq 3/(ns)$, we have
$\frac{\dfrac{\sqrt{\sigma^{2}+16\sigma^{2}\ln n}}{n}}{\sigma\sqrt{6c_{4}}\sqrt{\dfrac{s\ln(1+2n)}{n}}}\leq\frac{\sqrt{3\ln n}}{\sqrt{c_{4}ns\ln(1+2n)}}\leq 1,$
which implies $\frac{\sqrt{\sigma^{2}+16\sigma^{2}\ln n}}{n}\leq\sigma\sqrt{6c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}$. Consequently,
$\sup_{\|\bm{u}\|=1,\|\bm{v}\|=1}\Bigl|\frac{1}{n}\sum_{i=1}^{n}\bm{u}^{\top}\bm{Z}_{i}^{11}\bm{v}\varepsilon_{i}\Bigr|\leq 2\,\sigma\sqrt{6c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}\leq 5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}.$
This completes the proof. ∎
Proof of Theorem 1 Without loss of generality, we assume $\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\leq\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|$. Subsequently, it is easy to verify $\|\bm{\beta}_{1}^{\star}\|\leq\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|$.
Consider the level set $M=\{\bm{\beta}_{1}\in\mathbb{R}^{s}:\tilde{L}_{n}(\bm{\beta}_{1})+\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1})\leq\tilde{L}_{n}(\bm{\beta}_{1}^{\star})+\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star})\}$. It is clear that
$\inf_{\bm{\beta}_{1}\in\mathbb{R}^{s}}\tilde{L}_{n}(\bm{\beta}_{1})+\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1})=\inf_{\bm{\beta}_{1}\in M}\tilde{L}_{n}(\bm{\beta}_{1})+\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}).$
Since $\tilde{L}_{n}(\cdot)+\tilde{P}_{\lambda_{n}}(\cdot)$ is continuous and the level set is compact, there exists at least one minimizer $\bm{\hat{\beta}}_{1}$ in the level set.
Noting that condition (iv) of Assumption 1 implies that all sub-differential and derivatives of $P_{\lambda_{n}}$ are bounded in magnitude by $\lambda_{n}\varrho$, we get
$|\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})-\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star})|\leq\lambda_{n}\varrho\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|_{1}\leq\lambda_{n}\sqrt{s}\varrho\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|.$ (13)
According to the definition of $\bm{\hat{\beta}}_{1}^{o}$, for any $\bm{\beta}_{1}\in\mathbb{R}^{s}$,
$\tilde{L}_{n}(\bm{\hat{\beta}}_{1}^{o})+\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})\leq\tilde{L}_{n}(\bm{\beta}_{1})+\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}).$
Then we get
$0$ $\geq\tilde{L}_{n}(\bm{\hat{\beta}}_{1}^{o})+\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})-\tilde{L}_{n}(\bm{\beta}_{1}^{\star})-\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star})$
$=\frac{1}{4n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o\top}\bm{Z}_{i}^{11}\bm{\hat{\beta}}_{1}^{o}-y_{i})^{2}-\frac{1}{4n}\sum_{i=1}^{n}(\bm{\beta}_{1}^{\star\top}\bm{Z}_{i}^{11}\bm{\beta}_{1}^{\star}-y_{i})^{2}+\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})-\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star})$
$=\frac{1}{4n}\sum_{i=1}^{n}\bigl[(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})\bigr]^{2}$
$\quad-\frac{1}{2n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})\varepsilon_{i}$
$\quad+\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})-\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star})$
$\geq\frac{c_{3}}{4}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|^{2}\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|^{2}$
$\quad-\frac{1}{2n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})\varepsilon_{i}$
$\quad+\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})-\tilde{P}_{\lambda_{n}}(\bm{\beta}_{1}^{\star}).$
Combining the inequality (13), we obtain
$\frac{c_{3}}{4}\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|^{2}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|^{2}$
$\leq$ $\frac{1}{2n}\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|\sum_{i=1}^{n}(\frac{\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}}{\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|})^{\top}\bm{Z}_{i}^{11}(\frac{\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}}{\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|})\varepsilon_{i}$
$\qquad+\lambda_{n}\sqrt{s}\varrho\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|.$
Under the event $E^{1}$, we get
$\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\leq\frac{10\sigma\sqrt{c_{4}}}{c_{3}\|\bm{\beta}^{\star}\|}\sqrt{\frac{s\ln(1+2n)}{n}}+\frac{4\lambda_{n}\varrho\sqrt{s}}{c_{3}\|\bm{\beta}^{\star}\|^{2}}.$
which yields that
$\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\leq$ $\frac{10\sigma\sqrt{c_{4}}}{c_{1}c_{3}}\sqrt{\frac{\ln(1+2n)}{n}}+\frac{4\lambda_{n}\varrho}{c_{1}^{2}c_{3}\sqrt{s}}$
$<$ $C_{0}\big(\sqrt{\frac{\ln(1+2n)}{n}}+\frac{\lambda_{n}\varrho}{\sqrt{s}}\big),$ (14)
where $C_{0}=\max\left\{10\sigma\sqrt{c_{4}}/c_{1}c_{3},4/c_{1}^{2}c_{3}\right\}$. Combing this and Lemma 1, we get the desired result. ∎
To prove Theorem 2, we also need the following lemma.
Lemma 2.
Define
$E^{2}=:\Big\{\max_{1\leq j\leq d-s}\sup_{\|\bm{u}\|\leq r_{n}}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u}\varepsilon_{i}|\leq t_{2}+\frac{\sqrt{s(\sigma^{2}+16\sigma^{2}\ln n)}}{n}\Big\}$
with $t_{2}=\sqrt{2}\sigma r_{n}\sqrt{\frac{s(s+1)\ln(1+2n)+s\ln d}{n}}$. Under Assumptions 2-4, it follows that
$\mathbb{P}\big(E^{2}\big)\geq 1-\frac{1}{n}-\frac{1}{n^{-n/2}}.$
Proof.
Let $\mathbb{B}(\bm{u},\delta):=\left\{\bm{v}\in\mathbb{R}^{s}:\|\bm{v}-\bm{u}\|\leq\delta\right\}$. Then follows from the Lemma 14.27 of Bühlmann and Van De Geer (2011) that there exist $K:=(1+2n)^{s}$ vectors $\bm{u_{k}}$ with $\bm{u_{k}}\in\mathbb{R}^{s}$ and $\|\bm{u_{k}}\|\leq r_{n}$ such that
$\left\{\bm{u}\in\mathbb{R}^{s}:~~~\|\bm{u}\|\leq r_{n}\right\}\subseteq\cup_{k=1}^{K}\mathbb{B}(\bm{u_{k}},\frac{1}{n}).$
Using this fact, we obtain the following inequalities
$\begin{split}&\max_{1\leq j\leq d-s}\sup_{\|\bm{u}\|\leq r_{n}}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u}\varepsilon_{i}|\\ \leq&\max_{1\leq j\leq d-s}\sup_{\bm{u}\in\cup_{k=1}^{K}\mathbb{B}(\bm{u}_{k},\frac{1}{n})}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u}\varepsilon_{i}|\\ \leq&\max_{1\leq j\leq d-s}\max_{1\leq k\leq K}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u_{k}}\varepsilon_{i}|\\ &\qquad+\max_{1\leq j\leq d-s}\sup_{\bm{u}\in\mathbb{B}(\bm{u_{k}},\frac{1}{n})}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}(\bm{u}-\bm{u_{k}})\varepsilon_{i}|.\end{split}$ (15)
Noting the fact that $|\bm{\beta}^{\top}\bm{Z}\bm{\beta}^{\prime}|\leq|\bm{Z}|_{\infty}\|\bm{\beta}\|_{1}\|\bm{\beta}^{\prime}\|_{1}$ for any $n\times{}d$ matrix $\bm{Z}$ and vectors $\bm{\beta},\,\bm{\beta}^{\prime}\in\mathbb{R}^{d}$, one can conclude from (4) that
$\sum_{i=1}^{n}(\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u})^{2}\leq\sum_{i=1}^{n}\|\bm{e}_{d-s,j}\|_{1}^{2}|\bm{Z}_{i}^{21}|_{\infty}^{2}\|\bm{u}\|_{1}^{2}\leq\sum_{i=1}^{n}|\bm{Z}_{i}|_{\infty}^{2}\|\bm{u}\|_{1}^{2}\leq n\|\bm{u}\|_{1}^{2}$ (16)
for any $\bm{u}\in\mathbb{R}^{s}$.
For the first term of (15), we can apply the countable subadditivity of probability, Bernstein’s inequality, inequality (16) and
$\|\bm{u}_{k}\|_{1}\leq\sqrt{s}\|\bm{u}_{k}\|\leq\sqrt{s}r_{n}$
to obtain, for any $t>0$,
$\begin{split}&\mathbb{P}\Big(\max_{1\leq j\leq d-s}\max_{1\leq k\leq K}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u_{k}}\varepsilon_{i}|>t\Big)\\ \leq&\sum_{j=1}^{d-s}\sum_{k=1}^{K}\mathbb{P}\big(|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u_{k}}\varepsilon_{i}|>t\big)\\ \leq&2dK\exp(-\frac{n^{2}t^{2}}{2\sigma^{2}\sum_{i=1}^{n}(\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u_{k}})^{2}})\\ \leq&2dK\exp(-\frac{nt^{2}}{2\sigma^{2}sr_{n}^{2}})\\ \leq&2\exp\left\{-\frac{nt^{2}}{2\sigma^{2}sr_{n}^{2}}+s\ln(1+2n)+\ln d\right\}.\end{split}$
Taking $t=t_{2}$, we drive
$\begin{split}&\mathbb{P}\Big(\max_{1\leq j\leq d-s}\max_{1\leq k\leq K}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u_{k}}\varepsilon_{i}|>t_{2}\Big)<\frac{1}{n}.\end{split}$
For the second term of (15), we can calculate that
$\begin{split}&\max_{1\leq j\leq d-s}\sup_{\bm{u}\in\mathbb{B}(\bm{u_{k}},\frac{1}{n})}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}(\bm{u}-\bm{u_{k}})\varepsilon_{i}|\\ \leq&\max_{1\leq j\leq d-s}\sup_{\bm{u}\in\mathbb{B}(\bm{u_{k}},\frac{1}{n})}\sqrt{\frac{1}{n}\sum_{i=1}^{n}[\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}(\bm{u}-\bm{u_{k}})]^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}}\\ \leq&\max_{1\leq j\leq d-s}\sup_{\bm{u}\in\mathbb{B}(\bm{u_{k}},\frac{1}{n})}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\bm{e}_{d-s,j}\|_{1}^{2}|\bm{Z}_{i}^{21}|_{\infty}^{2}\|\bm{u}-\bm{u_{k}}\|_{1}^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}}\\ \leq&\sup_{\bm{u}\in\mathbb{B}(\bm{u_{k}},\frac{1}{n})}\sqrt{\frac{s}{n}\sum_{i=1}^{n}|\bm{Z}_{i}^{21}|_{\infty}^{2}\|\bm{u}-\bm{u_{k}}\|_{2}^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}}\\ \leq&\frac{\sqrt{s}}{n}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}}.\end{split}$ (17)
Combing this, inequalities (12), (15) and (17), get the desired result. ∎
Lemma 3.
Under Assumptions 2-4, it follows that
$\mathbb{P}\big(E^{3}\big)\geq 1-\frac{1}{n}$
where the event $E^{3}$ is defined by
$E^{3}:=\{\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star}\varepsilon_{i}|\leq\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}\}.$
Proof.
From Bernstein’s inequality, (16) and (3), we can conclude that
$\begin{split}\mathbb{P}\Big(\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star}\varepsilon_{i}|>t\Big)\leq&2d\exp\Big\{\frac{-n^{2}t^{2}}{2\sigma^{2}\sum_{i=1}^{n}(\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star})^{2}}\Big\}\\ \leq&2d\exp\Big\{\frac{-nt^{2}}{2\sigma^{2}\|\bm{\beta}_{1}^{\star}\|_{1}^{2}}\Big\}\\ \leq&2\exp\big\{\frac{-nt^{2}}{2\sigma^{2}c_{2}^{2}s^{2}}+\ln d\big\}.\end{split}$
By taking $t=t_{1}:=\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}$, we subsequently obtain
$\mathbb{P}\Big(\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star}\varepsilon_{i}|>t_{1}\Big)<\frac{1}{n}.$
Then, we get the desired result. ∎
By combining the ideas from the proofs of Theorem 3.3 in Chen et al. (2013) and Theorem 2.4 in Bai et al. (2024), we derive the following result.
Lemma 4.
Suppose Assumption 1 holds. Let $\bm{\beta}$ be a stationary point of Problem (2), i.e.,
$0\in\nabla L_{n}(\bm{\beta})+\partial P_{\lambda_{n}}(\bm{\beta}).$
If for any nonzero vector $\bm{g}\in\mathbb{R}^{d}$ with $g_{j}=0$ for all $j\notin\Gamma^{*}$,
$\bm{g}^{\top}\nabla^{2}L_{n}(\bm{\beta})\bm{g}+\sum_{j\in\Gamma^{*}}\min\partial^{2}p_{\lambda_{n}}(|\beta_{j}|)\,g_{j}^{2}>0,$ (18)
then $\bm{\beta}$ is a strict local minimum of Problem (2).
Proof of Theorem 2 According to Theorem 1, we can obtain the existence of a estimator $\bm{\hat{\beta}}_{1}^{o}\in\mathbb{R}^{s}$, and the error bound between $\bm{\hat{\beta}}_{1}^{o}$ and $\bm{\beta}_{1}^{\star}$ is $r_{n}$. Let $\bm{\hat{\beta}}$ be a vector with $\bm{\hat{\beta}}_{\Gamma^{\star}}=\bm{\hat{\beta}}_{1}^{o}$ and $\bm{\hat{\beta}}_{\Gamma^{\star c}}=0$.
Firstly, we prove that $\bm{\hat{\beta}}$ is a stationary point of Problem (2), i.e.,
[…]
By simple caiculation, we have
[…]
Since $\nabla\tilde{L}_{n}(\bm{\hat{\beta}}_{1}^{o})+\nabla\tilde{P}_{\lambda_{n}}(\bm{\hat{\beta}}_{1}^{o})=0,$ we only need to prove
$0\in\left[\nabla L_{n}(\bm{\hat{\beta}})+\partial P_{\lambda_{n}}(\bm{\hat{\beta}})\right]_{\Gamma^{\star c}}.$
It suffices to prove that
$\|\nabla L_{n}(\bm{\hat{\beta}})_{\Gamma^{\star c}}\|_{\infty}\leq\lambda_{n}\varrho,$ (20)
since $\partial p_{\lambda_{n}}(|\hat{\beta}_{j}|)\in[-\lambda_{n}\varrho,\lambda_{n}\varrho]$. Noting
$\begin{split}\nabla L_{n}(\bm{\hat{\beta}})_{\Gamma^{\star c}}&=\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}^{\top}\bm{Z}_{i}\bm{\hat{\beta}}-y_{i})\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\\ &=\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})Z_{i}^{21}\bm{\hat{\beta}}_{1}^{o}-\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\varepsilon_{i},\end{split}$
we get
$\begin{aligned} &\|\nabla L_{n}(\bm{\hat{\beta}})_{\Gamma^{\star c}}\|_{\infty}\\ \leq&\|\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\|_{\infty}+\|\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\varepsilon_{i}\|_{\infty}.\end{aligned}$ (21)
We first estimate the upper bound of the first term on the right-hand side of the above inequality. Using again the fact that $|\bm{\beta}^{\top}\bm{Z}\bm{\beta}^{\prime}|\leq|\bm{Z}|_{\infty}\|\bm{\beta}\|_{1}\|\bm{\beta}^{\prime}\|_{1}$ for any $n\times{}d$ matrix $\bm{Z}$ and vectors $\bm{\beta},\,\bm{\beta}^{\prime}\in\mathbb{R}^{d}$, we conclude from Hölder’s inequality, Assumption 4 and (3) that
$\begin{aligned} &\|\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\|_{\infty}\\ =&\max_{1\leq k\leq d-s}\big|(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})^{\top}\big(\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})\bm{e}_{d-s,k}^{\top}\bm{Z}_{i}^{21}\big)\bm{\hat{\beta}}_{1}^{o}\big|\\ \leq&\max_{1\leq k\leq d-s}\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|_{1}\|\bm{\hat{\beta}}_{1}^{o}\|_{1}\big|\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})\bm{e}_{d-s,k}^{\top}\bm{Z}_{i}^{21}\big|_{\infty}\\ =&\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|_{1}\|\bm{\hat{\beta}}_{1}^{o}\|_{1}\max_{1\leq k\leq d-s}\max_{1\leq j\leq s}\max_{1\leq l\leq s}\big|\frac{1}{n}\sum_{i=1}^{n}\bm{(}\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}\bm{e}_{s,j}\bm{e}_{d-s,k}^{\top}\bm{Z}_{i}^{21}\bm{e}_{s,l}\big|\\ \leq&\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|_{1}|\bm{\hat{\beta}}_{1}^{o}\|_{1}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|_{1}\max_{1\leq k\leq d-s}\max_{1\leq j\leq s}\big|\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{11}\bm{e}_{s,j}\bm{e}_{d-s,k}^{\top}\bm{Z}_{i}^{21}\big|_{\infty}\end{aligned}$
$\begin{aligned} \leq&s^{3/2}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}\||\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{11}\otimes\bm{Z}_{i}^{21}|_{\infty}\\ \leq&s^{3/2}\left(\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|+2\|\bm{\beta}_{1}^{\star}\|\right)\left(\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|+\|\bm{\beta}_{1}^{\star}\|\right)\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\frac{c_{5}}{\sqrt{n}}\\ \leq&6s^{3/2}\|\bm{\beta}_{1}^{\star}\|^{2}r_{n}\frac{c_{5}}{\sqrt{n}}\\ \leq&6c_{2}^{2}s^{5/2}r_{n}\frac{c_{5}}{\sqrt{n}}\end{aligned}$ (22)
[…]
$r_{n}<\frac{c_{3}c_{1}^{2}\sqrt{s}}{2(4c_{3}+3c_{4})c_{2}}<c_{1}\sqrt{s}\leq|\bm{\beta}_{1}^{\star}|.$
We proceed to estimate the upper bound of the second term on the right-hand side of inequality (21). Notice that
$\begin{split}&\|\frac{1}{n}\sum_{i=1}^{n}\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\varepsilon_{i}\|_{\infty}\\ =&\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\hat{\beta}}_{1}^{o}\varepsilon_{i}|\\ \leq&\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})\varepsilon_{i}|+\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star}\varepsilon_{i}|\end{split}$
$\begin{split}\leq&\max_{1\leq j\leq d-s}\sup_{\|\bm{u}\|\leq r_{n}}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{u}\varepsilon_{i}|+\max_{1\leq j\leq d-s}|\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{d-s,j}^{\top}\bm{Z}_{i}^{21}\bm{\beta}_{1}^{\star}\varepsilon_{i}|\\ \leq&t_{2}+\frac{\sqrt{s(\sigma^{2}+t_{2})}}{n}+\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}.\end{split}$
under the event $E^{2}\cap E^{3}$. Therefore, combining the above inequality, inequalities (21) and (22), we get
$\begin{aligned} &\|\nabla L_{n}(\bm{\hat{\beta}})_{\Gamma^{\star c}}\|_{\infty}\\ \leq&6c_{2}^{2}s^{5/2}r_{n}\frac{c_{5}}{\sqrt{n}}+t_{2}+\frac{\sqrt{s(\sigma^{2}+16\sigma^{2}\ln n)}}{n}+\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}\end{aligned}$ (23)
under the event $E^{1}\cap E^{2}\cap E^{3}$.
On the other hand, note that the left inequality of condition (7) results in
$\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}\leq\frac{1}{6}\lambda_{n}\varrho.$ (24)
By the first inequality of condition (5), we get
$6c_{2}^{2}c_{5}C_{0}\frac{s^{5/2}}{\sqrt{n}}\frac{\lambda_{n}\varrho}{\sqrt{s}}\leq\frac{1}{6}\lambda_{n}\varrho$
which together with the left inequality of condition (7) also leads to
$\begin{split}6c_{2}^{2}c_{5}C_{0}\frac{s^{5/2}}{\sqrt{n}}\sqrt{\frac{\ln(1+2n)}{n}}=&\frac{6c_{2}^{2}c_{5}C_{0}s^{2}}{\sqrt{n}}\sqrt{\frac{s\ln(1+2n)}{n}}\\ \leq&\frac{1}{6}\sqrt{\frac{s\ln(1+2n)}{n}}\\ \leq&\frac{1}{6}\lambda_{n}\varrho.\end{split}$
By simple calculation, we get
$\begin{split}&\frac{\sqrt{2}C_{0}\sigma\sqrt{\frac{s(s+1)\ln(1+2n)+s\ln d}{n}}\sqrt{\frac{\ln(1+2n)}{n}}}{\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}}\\ =&\frac{C_{0}}{c_{2}}\sqrt{\frac{\ln(1+2n)}{n}}\sqrt{\frac{1}{s}+\frac{\ln(1+2n)}{\ln(1+2n)+\ln d}}\\ <&\frac{C_{0}}{c_{2}}\sqrt{\frac{\ln(1+2n)}{n}}\sqrt{2}\\ \leq&1\end{split}$
where the last ineauality follows from the second inequality of condition (5). Combing this and the inequality (24) enable us to get
$\sqrt{2}C_{0}\sigma\sqrt{\frac{s(s+1)\ln(1+2n)+s\ln d}{n}}\sqrt{\frac{\ln(1+2n)}{n}}\leq\frac{1}{6}\lambda_{n}\varrho.$
By simple calculation, we get
$\begin{split}\frac{\sqrt{2}C_{0}\sigma\sqrt{\frac{s(s+1)\ln(1+2n)+s\ln d}{n}}\frac{\lambda_{n}\varrho}{\sqrt{s}}}{\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}}=&\frac{\lambda_{n}\varrho C_{0}}{c_{2}\sqrt{s}}\sqrt{\frac{1}{s}+\frac{\ln(1+2n)}{\ln(1+2n)+\ln d}}\\ <&\frac{\sqrt{2}\lambda_{n}\varrho C_{0}}{c_{2}\sqrt{s}}\leq 1\end{split}$
where the last ineauality follows from the second inequality of condition (6). Combing this and the inequality (24) enable us to get
$\sqrt{2}C_{0}\sigma\sqrt{\frac{s(s+1)\ln(1+2n)+s\ln d}{n}}\times\frac{\lambda_{n}\varrho}{\sqrt{s}}\leq\frac{1}{6}\lambda_{n}\varrho.$
Considering
$\begin{split}\frac{\frac{\sqrt{s(\sigma^{2}+16\sigma^{2}\ln n)}}{n}}{\sqrt{2}c_{2}\sigma s\sqrt{\frac{\ln(1+2n)+\ln d}{n}}}=&\frac{1}{c_{2}\sqrt{ns}}\times\frac{\sqrt{\frac{1}{2}+8\ln n}}{\sqrt{\ln(1+2n)+\ln d}}\\ <&\frac{3}{c_{2}\sqrt{ns}}\leq 1\end{split}$
where first inequality derives from $n\geq 3$(resulting in $\ln n>1/2$) and $\ln n\leq\ln(1+2n)+\ln d$ and the last ineauality follows from the second inequality of condition (6). Combing this and the inequality (24) enable us to get
$\frac{\sqrt{s(\sigma^{2}+16\sigma^{2}\ln n)}}{n}\leq\frac{1}{6}\lambda_{n}\varrho.$
By these inequalities, the definition of $r_{n}$ and inequality (23), we get (20) under the event $E^{1}\cap E^{2}\cap E^{3}$.
Next, we prove that $\bm{\hat{\beta}}$ is a local minimizer if the event $E^{1}\cap E^{2}\cap E^{3}$ occurs. Based on Lemma 4, it suffices to prove that $\bm{\hat{\beta}}$ satisfies condition (18). Note that the convexity of $p_{\lambda_{n}}(t)+\frac{\mu}{2}t^{2}$ implies
$\min\partial^{2}p_{\lambda}(t)+\mu\geq 0~~\mbox{for any}~~t\geq 0,$
and hence
$\sum_{j\in\Gamma^{*}}(\min\partial^{2}p_{\lambda_{n}}(|\hat{\beta}_{j}|)+\mu)g_{j}^{2}\geq 0.$
On the other hand, a simple calculation yields that
$\begin{split}&\bm{g}^{\top}\nabla^{2}L_{n}(\bm{\hat{\beta}})\bm{g}+\sum_{j\in\Gamma^{\star}}\min\partial^{2}p_{\lambda_{n}}(|\bm{\hat{\beta}}_{j}|)g_{j}^{2}\\ =&\bm{g}^{T}\nabla^{2}L_{n}(\bm{\hat{\beta}})\bm{g}+\sum_{j\in\Gamma^{\star}}(\min\partial^{2}p_{\lambda_{n}}(|\bm{\hat{\beta}}_{j}|)+\mu)\bm{g}_{j}^{2}-\sum_{j\in\Gamma^{\star}}\mu g_{j}^{2}.\end{split}$
Therefore, it remains to prove that for any nonzero vector $\bm{g}\in\mathbb{R}^{d}$ with $g_{j}=0$ for all $j\notin\Gamma^{*}$,
$\bm{g}^{\top}\nabla^{2}L_{n}(\hat{\bm{\beta}})\bm{g}-\sum_{j\in\Gamma^{\star}}\mu g_{j}^{2}>0.$ (25)
We now proceed to prove this inequality. Since
$|\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}\varepsilon_{i}|\leq\|\bm{g}_{1}\|^{2}\sup_{\|\bm{u}\|=1,\|\bm{v}\|=1}|\frac{1}{n}\sum_{i=1}^{n}\bm{u}^{\top}\bm{Z}_{i}^{11}\bm{v}\varepsilon_{i}|,$
we have
$|\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}\varepsilon_{i}|\leq 5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}\|\bm{g}_{1}\|^{2}$
if the event $E^{1}$ occurs. Notice that
$\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})g_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}$
$\leq$ $\frac{1}{n}\sqrt{\sum_{i=1}^{n}\big((\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})\big)^{2}}\sqrt{\sum_{i=1}^{n}\big(g_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}\big)^{2}}.$
Combining these two inequalities and Assumption 3, under the event $E^{1}$, we have for any non-zero vector $\bm{g}\in\mathbb{R}^{d}$ with $g_{j}=0$ for all $j\notin\Gamma^{*}$,
$\begin{split}&\bm{g}^{\top}\nabla^{2}L_{n}(\bm{\hat{\beta}})\bm{g}-\sum_{j\in\Gamma^{\star}}\mu g_{j}^{2}\\ =&\frac{2}{n}\sum_{i=1}^{n}\bm{g}^{\top}\bm{Z}_{i}\bm{\hat{\beta}}\bm{\hat{\beta}}^{\top}\bm{Z}_{i}\bm{g}+\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}^{\top}\bm{Z}_{i}\bm{\hat{\beta}}-y_{i})\bm{g}^{\top}\bm{Z}_{i}\bm{g}-\sum_{j\in\Gamma^{\star}}\mu g_{j}^{2}\\ =&{\frac{2}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o\top}\bm{Z}_{i}^{11}\bm{g}_{1})^{2}+\frac{1}{n}\sum_{i=1}^{n}(\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star})^{\top}\bm{Z}_{i}^{11}(\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star})g_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}}\\ &\qquad-\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{1}^{\top}\bm{Z}_{i}^{11}\bm{g}_{1}\varepsilon_{i}-\sum_{j\in\Gamma^{\star}}\mu g_{j}^{2}\\ \geq&2c_{3}\|\bm{\hat{\beta}}_{1}^{o}\|^{2}\|\bm{g}_{1}\|^{2}-c_{4}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\|\bm{g}_{1}\|^{2}\\ &\qquad\qquad-5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}\|\bm{g}_{1}\|^{2}-\mu\|\bm{g}_{1}\|^{2}\\ \geq&2c_{3}\|\bm{\hat{\beta}}_{1}^{o}\|^{2}\|\bm{g}_{1}\|^{2}-c_{4}\|\bm{\hat{\beta}}_{1}^{o}+\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|\|\bm{g}_{1}\|^{2}-c_{3}\|\bm{\beta}_{1}^{\star}\|^{2}\|\bm{g}_{1}\|^{2}-\mu\|\bm{g}_{1}\|^{2}\\ \geq&\Big(2c_{3}(||\bm{\beta}_{1}^{\star}\|^{2}+\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|^{2}-2||\bm{\beta}_{1}^{\star}\|\|\bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|)\Big)\|g_{1}\|^{2}\\ &-\Big(c_{4}(2\|\bm{\beta}_{1}^{\star}\|+\|{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|)\|\ \bm{\hat{\beta}}_{1}^{o}-\bm{\beta}_{1}^{\star}\|+5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}+\mu\Big)\|\bm{g}_{1}\|^{2}\\ \geq&\Big(2c_{3}\|\bm{\beta}_{1}^{\star}\|^{2}-4c_{3}\|\bm{\beta}_{1}^{\star}\|r_{n}-c_{4}r_{n}^{2}-2c_{4}\|\bm{\beta}_{1}^{\star}\|r_{n}-5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}-\mu\Big)\|\bm{g}_{1}\|^{2}\\ \geq&\Big(2c_{3}c_{1}^{2}s-c_{4}r_{n}^{2}-(4c_{3}+2c_{4})c_{2}\sqrt{s}r_{n}-5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}-\mu\Big)\|\bm{g}_{1}\|^{2}.\end{split}$
$\begin{split}\geq&\Big(2c_{3}c_{1}^{2}s-(4c_{3}+3c_{4})c_{2}\sqrt{s}r_{n}-5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}-\mu\Big)\|\bm{g}_{1}\|^{2}\\ \geq&\frac{1}{2}\big(2c_{3}c_{1}^{2}s-(4c_{3}+3c_{4})c_{2}\sqrt{s}r_{n}-5\sigma\sqrt{c_{4}}\sqrt{\frac{s\ln(1+2n)}{n}}-\mu\big)\|\bm{g}_{1}\|^{2}.\end{split}$
where the fifth inequality follows from $r_{n}\leq c_{1}\sqrt{s}$ and $c_{1}\leq c_{2}$. Combing the above inequality and the condition $\mu\leq c_{1}^{2}c_{3}s$ and (6), we get (25) under the event $E^{1}$.
Overall, we prove that $\bm{\hat{\beta}}$ is a local minimizer under the event $E^{1}\cap E^{2}\cap E^{3}$. By Lemmas 1, 2 and 2, we complete the proof.∎
[…]
Proof of Proposition 1 Based on the similar proof method to that of Theorem 5 in Fan et al. (2018), it is easy to prove equation (3.6) and thus it is omitted. We mainly prove equation (3.7).
For any $\tau>0$, define the following auxiliary function
$G_{\tau}(\bm{\beta}|\bm{\gamma}):=L(\bm{\gamma})+\langle\nabla L(\bm{\gamma}),\bm{\beta}-\bm{\gamma}\rangle+\frac{1}{2\tau}\|\bm{\beta}-\bm{\gamma}\|^{2}+P_{\lambda}(\bm{\gamma})+\sum_{j=1}^{d}w_{\lambda,j}(|\beta_{j}|-|\gamma_{j}|)$ (1)
where
$w_{\lambda,j}=\begin{cases}\max(p_{\lambda}^{{}^{\prime}}(|\gamma_{j}|),\epsilon_{1}),&\text{if }\gamma_{j}\neq 0,\\ \lambda\varrho,&\text{if }\gamma_{j}=0.\end{cases}$
It is clear that minimizing (3) with respect to $\bm{\beta}$ is equivalent to the following minimization problem
$\min_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\bm{\beta}-(\bm{\gamma}+\tau\nabla L(\bm{\gamma}))\|^{2}+\tau\sum_{j=1}^{d}w_{\lambda,j}|\beta_{j}|.$ (2)
For any $r>0$, let $B_{r}=\{\bm{\beta}\in\mathbb{R}^{d}:\|\bm{\beta}\|\leq r\}$ and $\hat{L}=\sup_{\bm{\beta}\in B_{r}}\|\nabla^{2}L(\bm{\beta})\|.$ Then for any $\tau\in(0,{\hat{L}}^{-1}]$ and $\bm{\beta},\bm{\gamma}\in B_{r}$, we have
$\begin{split}F(\bm{\beta})&=L(\bm{\gamma})+\langle\nabla L(\bm{\gamma}),\bm{\beta}-\bm{\gamma}\rangle+\frac{1}{2}(\bm{\beta}-\bm{\gamma})^{\top}\nabla^{2}L(\bm{\xi})(\bm{\beta}-\bm{\gamma})+P_{\lambda}(\bm{\beta})\\ &\leq L(\bm{\gamma})+\langle\nabla L(\bm{\gamma}),\bm{\beta}-\bm{\gamma}\rangle+\frac{1}{2}(\bm{\beta}-\bm{\gamma})^{\top}\nabla^{2}L(\bm{\xi})(\bm{\beta}-\bm{\gamma})\\ &+P_{\lambda}(\bm{\gamma})+\sum_{j=1}^{n}v_{j}(|\beta_{j}|-|\gamma_{j}|)\\ &\leq L(\bm{\gamma})+\langle\nabla L(\bm{\gamma}),\bm{\beta}-\bm{\gamma}\rangle+\frac{1}{2}(\bm{\beta}-\bm{\gamma})^{\top}\nabla^{2}L(\bm{\xi})(\bm{\beta}-\bm{\gamma})\\ &+P_{\lambda}(\bm{\gamma})+\sum_{\gamma_{j}\neq 0}\max(p_{\lambda}^{{}^{\prime}}(|\gamma_{j}|),\epsilon_{1})(|\beta_{j}|-|\gamma_{j}|)+\sum_{\gamma_{j}=0}\lambda\varrho(|\beta_{j}|-|\gamma_{j}|)\end{split}$
$\begin{split}&\leq G_{\tau}(\bm{\beta}|\bm{\gamma})+\frac{1}{2}(\bm{\beta}-\bm{\gamma})^{\top}\nabla^{2}L(\bm{\xi})(\bm{\beta}-\bm{\gamma})-\frac{1}{2\tau}\|\bm{\beta}-\bm{\gamma}\|^{2}\\ &\leq G_{\tau}(\bm{\beta}|\bm{\gamma})+\frac{\hat{L}}{2}\|\bm{\beta}-\bm{\gamma}\|^{2}-\frac{1}{2\tau}\|\bm{\beta}-\bm{\gamma}\|^{2}\\ &\leq G_{\tau}(\bm{\beta}|\bm{\gamma}).\end{split}$
where $v_{j}\in\partial p_{\lambda}(|\gamma_{j}|)$ and $\bm{\xi}=\bm{\gamma}+\alpha(\bm{\beta}-\bm{\gamma})$ for some $\alpha\in(0,1)$. The first inequality follows from the concavity of $p_{\lambda}$, the second inequality follows from the upper bound of $\partial p_{\lambda}$ and the third inequality follows from $\|\bm{\xi}\|\leq r$.
[…]
It is clear that problem (2) can be solved using the soft-thresholding operator $\mathcal{S}$, therefore we can get (10). ∎
To prove Proposition 2, we also need the following lemma.
Lemma 5.
Let $g_{k}=\|\nabla{}\ell(\beta^{k})\|_{2}$, $G_{k}=\sup_{\beta\in{}B_{k}}\|\nabla^{2}\ell(\beta)\|_{2}$ where $B_{k}=\{\beta\in\mathbb{R}^{p}:\|\beta\|_{2}\leq\|\beta^{k}\|_{2}+g_{k}\}.$ For any $\delta>0,\alpha_{0},\gamma_{1}\in(0,1)$, define
$j_{k}=\left\{\begin{array}[]{lll}0,&\mbox{if}~\gamma_{1}(G_{k}+\delta)\leq 1;\\ -\mbox{[}\log_{\gamma_{0}}\gamma_{1}(G_{k}+\delta)\mbox{]}+1,&\mbox{otherwise}.\end{array}\right$.
Then (11) holds, and there exists a nonnegative integer $\bar{j}$ such that $\tau_{k}\in[\gamma_{1}\gamma_{0}^{\bar{j}},\gamma_{1}].$
Proof sketch The proof follows the same structure as that of Lemmas 8 and 9 in Fan et al. (2018). The core function $L$ is identical. Although the regularizer considered here (a weakly convex concave regularizer or a weighted $\ell_{1}$ penalty) differs in form from the $\ell_{q}(0<q<1)$ regularizer in Fan et al. (2018), the argument relies only on the coercivity of the regularizer and the nonexpansiveness of the proximal operator, both of which hold in our setting. Consequently, the entire line of reasoning applies directly, and detailed calculations are omitted. ∎
Proof of Proposition 2 Regarding Algorithm 1, based on the similar proof method to that of Theorem S3.1 in Fan et al. (2018), it is easy to prove it and so is omitted. The conclusion for Algorithm 2 is proved below, where the proof of (i) follows the approach of Proposition 3.2 in Bai et al. (2024). For any $\tau>0$, define the following auxiliary function
$G^{k}_{\tau}(\bm{\beta},\bm{\beta}^{k}):=L(\bm{\beta}^{k})+\langle\nabla L(\bm{\beta}^{k}),\bm{\beta}-\bm{\beta}^{k}\rangle+\frac{1}{2\tau}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}+P_{\lambda}(\bm{\gamma})+\sum_{j=1}^{d}w_{\lambda,j}^{k}(|\beta_{j}|-|\beta_{j}^{k}|).$ (3)
where
$w_{\lambda,j}^{k}=\begin{cases}\max(p_{\lambda}^{{}^{\prime}}(|\beta_{j}^{k}|),\epsilon_{1}),&\text{if }\beta_{j}^{k}\neq 0,\\ \lambda\varrho,&\text{if }\beta_{j}^{k}=0.\end{cases}$
Notice that the iteration $\bm{\beta}^{k+1}=\mathcal{S}(\bm{\beta}^{k}-\tau_{k}\nabla L(\bm{\beta}^{k}),{\lambda}\tau_{k}\bm{w}^{k})$ is equivalent to
$\bm{\beta}^{k+1}=\arg\min\frac{1}{2}\|\bm{\beta}-\big(\bm{\beta}^{k}-\tau_{k}\nabla L(\bm{\beta}^{k})\big)\|^{2}+\lambda\tau_{k}\|\bm{w}^{k}\circ\bm{\beta}\|_{1}$
Hence the sequence $\{\bm{\beta}^{k}\}$ is bounded because every entry of $\bm{w}^{k}$ is positive, $\lambda>0$ and $\tau_{k}\in[\gamma_{1}\gamma_{0}^{\bar{j}},\gamma_{1}]$ from Lemma 5. Then there exists a positive constant $G$ such that $G=\sup\|\nabla^{2}\ell(\bm{\beta}^{k})\|_{2}$.
(i) From the convavity of $p_{\lambda}^{{}^{\prime}}(t)>0$ for any $t>0$, we conclude that
$\begin{split}&\sum_{j=1}^{d}p_{\lambda}(|\beta_{j}^{k+1}|)-\sum_{j=1}^{d}p_{\lambda}(|\beta_{j}^{k}|)\\ \leq&\sum_{\beta_{j}^{k}\neq 0}p_{\lambda}^{{}^{\prime}}(|\beta_{j}^{k}|)(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)+\sum_{\beta_{j}^{k}=0}\xi_{j}^{k}(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)\end{split}$
$\begin{split}\leq&\sum_{\beta_{j}^{k}\neq 0}p_{\lambda}^{{}^{\prime}}(|\beta_{j}^{k}|)(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)+\sum_{\beta_{j}^{k}=0}\lambda\varrho(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)\\ =&\sum_{j=1}^{d}w_{\lambda,j}^{k}(|\beta_{j}^{k+1}|-|\beta_{j}^{k}|)\end{split}$ (4)
where the second inequality follows from $\lim\limits_{t\rightarrow 0^{+}}p_{\lambda}^{{}^{\prime}}(t)=\lambda\varrho$ and $\varrho>0$. Additionally, from the Taylor expansion, it follows that
$L(\bm{\beta}^{k})-L(\bm{\beta}^{k+1})\geq\langle\nabla L(\bm{\beta}^{k}),\bm{\beta}^{k}-\bm{\beta}^{k+1}\rangle-\frac{G}{2}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}.$ (5)
Combining (4) and (5), we can get
$\begin{split}&F(\bm{\beta}^{k})-F(\bm{\beta}^{k+1})\\ =&L(\bm{\beta}^{k})+\sum_{j=1}^{d}p_{\lambda}(|\beta_{j}^{k}|)-L(\bm{\beta}^{k+1})-\sum_{j=1}^{d}p_{{\lambda}}(|\beta_{j}^{k+1}|)\\ \geq&\langle\nabla L(\bm{\beta}^{k}),\bm{\beta}^{k}-\bm{\beta}^{k+1}\rangle-\frac{\hat{L}}{2}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}+\sum_{j=1}^{d}w_{\lambda,j}^{k}(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)\end{split}$ (6)
For each subproblem, $\bm{\beta}^{k+1}$ staisfies the optimal condition
$\bm{0}\in\nabla L(\bm{\beta}^{k})+\frac{1}{\tau_{k}}(\bm{\beta}^{k+1}-\bm{\beta}^{k})+\bm{w}_{\lambda}^{k}\circ\partial\|\bm{\beta}^{k+1}\|_{1},$
which implies that there exists an vector $\bm{\zeta}^{k+1}=(\zeta_{1}^{k+1},\cdots,\zeta_{d}^{k+1})^{\top}\in\partial\|\bm{\beta}^{k+1}\|_{1}$ such that
$\nabla L(\bm{\beta}^{k})+\frac{1}{\tau_{k}}(\bm{\beta}^{k+1}-\bm{\beta}^{k})+\bm{w}_{\lambda}^{k}\circ\bm{\zeta}^{k+1}=\bm{0},$ (7)
where $\bm{w}_{\lambda}^{k}=(w_{\lambda,1}^{k},\cdots,w_{\lambda,d}^{k})^{\top}$. Therefore, we obtain
$\begin{aligned} &G^{k}_{\tau_{k}}(\bm{\beta}^{k},\bm{\beta}^{k})-G^{k}_{\tau_{k}}(\bm{\beta}^{k+1},\bm{\beta}^{k})\\ =&\langle\nabla L(\bm{\beta}^{k}),\bm{\beta}^{k}-\bm{\beta}^{k+1}\rangle-\frac{1}{2\tau_{k}}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}+\sum_{j=1}^{d}w_{j}^{k}(|\beta_{j}^{k}|-|\beta_{j}^{k+1}|)\\ \geq&\langle\nabla L(\bm{\beta}^{k}),\bm{\beta}^{k}-\bm{\beta}^{k+1}\rangle+\sum_{j=1}^{d}\zeta_{j}^{k+1}w_{j}^{k}(\beta_{j}^{k}-\beta_{j}^{k+1})-\frac{1}{2\tau_{k}}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}\\ =&[\nabla L(\bm{\beta}^{k})+\frac{1}{\tau}(\bm{\beta}^{k+1}-\bm{\beta}^{k})+\sum_{j=1}^{d}w_{j}^{k}\zeta_{j}^{k+1}]^{\top}(\bm{\beta}^{k}-\bm{\beta}^{k+1})+\frac{1}{2\tau_{k}}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}\\ =&\frac{1}{2\tau_{k}}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2},\end{aligned}$
where the inequality follows from the convexity of $|\cdot|$ and $w_{\lambda,j}^{k}\geq 0$, and the last equality derives from (7). Combining this and (6), we get
$\begin{aligned} &F(\bm{\beta}^{k})-F(\bm{\beta}^{k+1})\\ \geq&G_{\tau_{k}}^{k}(\bm{\beta}^{k},\bm{\beta}^{k})-G_{\tau_{k}}^{k}(\bm{\beta}^{k+1},\bm{\beta}^{k})-\frac{G}{2}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}+\frac{1}{2\tau_{k}}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}\\ \geq&\left(\frac{1}{\tau_{k}}-\frac{G}{2}\right)\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}\\ \geq&0.\end{aligned}$ (8)
(ii) From the definition of $\bm{\beta}^{k+1}$ and (8), we have
$\sum_{k=0}^{K}[F(\bm{\beta}^{k})-F(\bm{\beta}^{k+1})]=F(\bm{\beta}^{0})-F(\bm{\beta}^{K+1})\geq\left(\frac{1}{\tau_{k}}-\frac{G}{2}\right)\sum_{k=0}^{K}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}.$
Hence, $\sum_{k=0}^{\infty}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}<\infty$, which implies $\lim\limits_{k\rightarrow\infty}\|\bm{\beta}^{k}-\bm{\beta}^{k+1}\|^{2}=0$.
(iii) Because $\{\bm{\beta}^{k}\}$ and $\{\tau_{k}\}$ have convergent subsequences, without loss of generality, assume that
$\bm{\beta}^{k}\to\bm{\beta}^{\diamond}\quad\text{and}\quad\tau_{k}\to\tau^{\diamond}\quad\text{as }k\to\infty,$ (9)
and denote
$G_{\tau}^{\diamond}(\bm{\beta},\bm{\beta}^{\diamond})=L(\bm{\beta}^{\diamond})+\langle\nabla L(\bm{\beta}^{\diamond}),\bm{\beta}-\bm{\beta}^{\diamond}\rangle+\frac{1}{2\tau}\|\bm{\beta}-\bm{\beta}^{\diamond}\|^{2}+P_{\lambda}(\bm{\gamma})+\sum_{j=1}^{d}w_{j}^{\diamond}(|\beta_{j}|-|\beta_{j}^{\diamond}|)$
with
$w_{j}^{\diamond}=\begin{cases}\max(p_{\lambda}^{{}^{\prime}}(|\beta_{j}^{\diamond}|),\epsilon_{1}),&\text{if }\beta_{j}^{\diamond}\neq 0,\\ \lambda\varrho,&\text{if }\beta_{j}^{\diamond}=0.\end{cases}$
It is known that the iteration $\bm{\beta}^{k+1}=\mathcal{S}(\bm{\beta}^{k}-\tau_{k}\nabla L(\bm{\beta}^{k}),{\lambda}\tau_{k}\bm{w}^{k}).$ Then for any $\bm{\beta}\in\mathbb{R}^{d}$,
$G^{k}_{\tau_{k}}(\bm{\beta}^{k+1},\bm{\beta}^{k})\leq G^{k}_{\tau_{k}}(\bm{\beta},\bm{\beta}^{k}).$
Combing the continuousness of the function $G^{k}_{\tau}(\cdot,\bm{\beta}^{k})$ and the limit (9), we can obtain
$G_{\tau^{\diamond}}^{\diamond}(\bm{\beta}^{\diamond},\bm{\beta}^{\diamond})=\lim_{k\to\infty}G_{\tau_{k}}^{k}(\bm{\beta}^{k+1},\bm{\beta}^{k})\leq\lim_{k\to\infty}G_{\tau_{k}}^{k}(\bm{\beta},\bm{\beta}^{k})=G_{\tau^{\diamond}}^{\diamond}(\bm{\beta},\bm{\beta}^{\diamond}).$
Based on the above inequality, we get
$\bm{\beta}^{\diamond}\in\arg\min_{\bm{\beta}\in\mathbb{R}^{d}}G_{\tau^{\diamond}}^{\diamond}(\bm{\beta},\bm{\beta}^{\diamond}).$
According to the Proposition 1, it means that $\bm{\beta}^{\diamond}$ is also satisfies
$\bm{\beta}=\mathcal{S}(\bm{\beta}-\tau\nabla L(\bm{\beta}),\tau\bm{w}_{\lambda}).$
Then, we complete the proof. ∎
[…]
Table 5: MSE(SD) of relative error for $d=128,s=10$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.091 (0.1414) 1.114 (0.1394) 1.091 (0.1646) 1.111 (0.1571) 1.152 (0.1743) 1.166 (0.1704) 1.187 (0.1779)
0.2 MSE 1.068 (0.1778) 1.108 (0.1568) 1.042 (0.1955) 1.03 (0.2445) 1.12 (0.3158) 1.133 (0.2951) 1.122 (0.349)
0.3 MSE 0.9705 (0.293) 1.052 (0.2643) 0.9269 (0.2784) 0.8665 (0.3783) 0.7603 (0.6056) 0.9416 (0.4998) 0.823 (0.5875)
0.4 MSE 0.5529 (0.5201) 0.7402 (0.4788) 0.6218 (0.407) 0.4724 (0.4664) 0.4097 (0.5788) 0.5031 (0.5934) 0.4627 (0.6022)
0.5 MSE 0.3079 (0.5235) 0.4849 (0.5386) 0.4257 (0.4788) 0.3526 (0.477) 0.2461 (0.4983) 0.2959 (0.5359) 0.2387 (0.4991)
0.6 MSE 0.05125 (0.2487) 0.2159 (0.4243) 0.2071 (0.3718) 0.06997 (0.2575) 0.04582 (0.2242) 0.08147 (0.2997) 0.04663 (0.2289)
0.7 MSE 0.07634 (0.302) 0.08627 (0.316) 0.1135 (0.3098) 0.05085 (0.2215) 0.07238 (0.2859) 0.07605 (0.3014) 0.05124 (0.2502)
0.8 MSE 0.01417 (0.1356) 0.008939 (0.08315) 0.02032 (0.141) 0.01137 (0.1108) 0.0005478 (0.0003627) 0.0003913 (0.0003058) 0.0004205 (0.0003104)
0.9 MSE 0.01311 (0.1256) 0.02341 (0.1631) 0.01126 (0.1101) 0.02173 (0.1514) 0.01205 (0.1157) 0.01312 (0.1279) 0.012 (0.1164)
1.0 MSE 0.0005213 (0.0002837) 0.0004813 (0.0008072) 0.008467 (0.08181) 0.0002522 (0.0001432) 0.0004765 (0.0002836) 0.0003329 (0.0002278) 0.0003577 (0.0002334)
Table 6: MSE(SD) of relative error for $d=128,s=10$ and $\sigma=0.05$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.085 (0.1328) 1.106 (0.1324) 1.087 (0.1423) 1.101 (0.1628) 1.153 (0.1718) 1.16 (0.1624) 1.175 (0.1777)
0.2 MSE 1.049 (0.1958) 1.101 (0.1642) 1.021 (0.2048) 1.043 (0.2243) 1.094 (0.3356) 1.121 (0.2632) 1.13 (0.342)
0.3 MSE 0.8273 (0.3511) 0.9234 (0.3213) 0.8148 (0.3337) 0.7896 (0.3815) 0.6387 (0.593) 0.6949 (0.5856) 0.6682 (0.612)
0.4 MSE 0.6436 (0.5085) 0.8282 (0.4484) 0.6899 (0.395) 0.5693 (0.4823) 0.4613 (0.5924) 0.59 (0.5957) 0.5024 (0.607)
0.5 MSE 0.243 (0.4811) 0.509 (0.5236) 0.4271 (0.471) 0.2844 (0.4427) 0.2429 (0.4918) 0.2886 (0.5111) 0.205 (0.4681)
0.6 MSE 0.203 (0.4607) 0.2593 (0.493) 0.2211 (0.405) 0.1493 (0.3698) 0.174 (0.4275) 0.1916 (0.454) 0.1882 (0.4472)
0.7 MSE 0.02853 (0.1732) 0.07949 (0.2845) 0.08671 (0.2745) 0.02283 (0.1482) 0.02904 (0.1793) 0.03935 (0.211) 0.04164 (0.2227)
0.8 MSE 0.003159 (0.002873) 0.03442 (0.1868) 0.05568 (0.2171) 0.02166 (0.1408) 0.002817 (0.002608) 0.0135 (0.1145) 0.01544 (0.1327)
0.9 MSE 0.0153 (0.1272) 0.01504 (0.1357) 0.01208 (0.1083) 0.01222 (0.1099) 0.01495 (0.1276) 0.01413 (0.1258) 0.01482 (0.1317)
1.0 MSE 0.002603 (0.001302) 0.0154 (0.1383) 0.01241 (0.1108) 0.01251 (0.112) 0.01539 (0.1304) 0.01415 (0.1254) 0.001726 (0.0009298)
Table 7: MSE(SD) of relative error for $d=256,s=15$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.121 (0.06421) 1.143 (0.05974) 1.138 (0.09982) 1.144 (0.07845) 1.18 (0.1058) 1.231 (0.08987) 1.263 (0.1073)
0.2 MSE 1.103 (0.15) 1.147 (0.1259) 1.062 (0.2094) 1.049 (0.2355) 1.087 (0.3919) 1.165 (0.3045) 1.15 (0.377)
0.3 MSE 0.9678 (0.3185) 1.101 (0.2053) 0.8716 (0.3311) 0.866 (0.3872) 0.7233 (0.6405) 0.8856 (0.5504) 0.7427 (0.6431)
0.4 MSE 0.5304 (0.5944) 0.8253 (0.482) 0.5715 (0.497) 0.4767 (0.5222) 0.3911 (0.5887) 0.5146 (0.6151) 0.4587 (0.6179)
0.5 MSE 0.2341 (0.476) 0.4294 (0.534) 0.293 (0.4405) 0.1878 (0.4069) 0.1399 (0.3996) 0.2263 (0.4728) 0.1496 (0.4082)
0.6 MSE 0.1276 (0.3851) 0.1752 (0.4071) 0.1024 (0.3097) 0.08888 (0.3035) 0.07372 (0.2932) 0.08987 (0.3296) 0.08984 (0.329)
0.7 MSE 0.06553 (0.2857) 0.08276 (0.3045) 0.04321 (0.2121) 0.04276 (0.21) 0.06264 (0.2731) 0.06505 (0.2842) 0.05078 (0.2491)
0.8 MSE 0.0003446 (0.000144) 0.02425 (0.1695) 0.01038 (0.1022) 0.02158 (0.1507) 0.0003082 (0.0001376) 0.0002127 (0.0001035) 0.0002275 (0.0001077)
0.9 MSE 0.0002978 (0.0001248) 0.0001692 (0.0001473) 0.01013 (0.09999) 0.0001356 (6.212e-05) 0.0002639 (0.0001202) 0.0001809 (9.108e-05) 0.0001931 (9.303e-05)
1.0 MSE 0.0002786 (0.0001441) 0.0001873 (0.0002725) 0.0001275 (6.871e-05) 0.0001276 (7.027e-05) 0.0002492 (0.0001398) 0.0001706 (0.0001048) 0.0001842 (0.0001158)
Table 8: MSE(SD) of relative error for $d=256,s=15$ and $\sigma=0.05$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.129 (0.07115) 1.151 (0.06742) 1.153 (0.07808) 1.175 (0.09573) 1.192 (0.1179) 1.245 (0.09804) 1.273 (0.1223)
0.2 MSE 1.12 (0.1201) 1.169 (0.09863) 1.114 (0.1472) 1.094 (0.2043) 1.221 (0.2625) 1.227 (0.2284) 1.219 (0.3407)
0.3 MSE 0.9922 (0.2916) 1.091 (0.2339) 0.9472 (0.2679) 0.9405 (0.3577) 0.8092 (0.6053) 0.95 (0.5133) 0.8946 (0.5978)
0.4 MSE 0.4294 (0.5732) 0.8253 (0.4619) 0.5233 (0.4976) 0.431 (0.5097) 0.3336 (0.5659) 0.4446 (0.6011) 0.3967 (0.6094)
0.5 MSE 0.2368 (0.4876) 0.5002 (0.5624) 0.2478 (0.4343) 0.1508 (0.3751) 0.1756 (0.4334) 0.2441 (0.4929) 0.2194 (0.4847)
0.6 MSE 0.05631 (0.2677) 0.1342 (0.3705) 0.1246 (0.3398) 0.05633 (0.2435) 0.06604 (0.2818) 0.07899 (0.3117) 0.06739 (0.29)
0.7 MSE 0.01509 (0.1325) 0.01446 (0.1328) 0.01168 (0.1076) 0.02301 (0.156) 0.01473 (0.1312) 0.02736 (0.1845) 0.01441 (0.132)
0.8 MSE 0.001777 (0.0008609) 0.01031 (0.09262) 0.01142 (0.106) 0.000823 (0.0004187) 0.001569 (0.0008123) 0.00113 (0.0006776) 0.00119 (0.0006817)
0.9 MSE 0.001553 (0.0006901) 0.0008882 (0.0005476) 0.01075 (0.09993) 0.000747 (0.0003353) 0.001367 (0.0006242) 0.0009917 (0.0004932) 0.001044 (0.0004984)
1.0 MSE 0.00136 (0.0005866) 0.0007221 (0.0004205) 0.0006343 (0.0002849) 0.0006249 (0.0002817) 0.0012 (0.0005614) 0.0008326 (0.0004163) 0.0008955 (0.0004457)
Table 9: Mean Time (SD) for $d=128,s=10$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 Time 0.03377 (0.01813) 0.1205 (0.09658) 0.01371 (0.007892) 0.01626 (0.01263) 0.1419 (0.09485) 0.3334 (0.1808) 0.55 (0.3274)
0.2 Time 0.02298 (0.007626) 0.05555 (0.02541) 0.003957 (0.00106) 0.005396 (0.009777) 0.06276 (0.03421) 0.08763 (0.01004) 0.2456 (0.1286)
0.3 Time 0.07842 (0.04432) 0.0535 (0.05498) 0.006069 (0.003454) 0.00785 (0.007068) 0.07034 (0.07559) 0.1158 (0.06931) 0.519 (0.3738)
0.4 Time 0.341 (0.1211) 0.1132 (0.08343) 0.02539 (0.01485) 0.04185 (0.04007) 0.1626 (0.1324) 0.2637 (0.2602) 1.015 (1.207)
0.5 Time 0.6451 (0.1132) 0.1928 (0.1897) 0.0556 (0.02597) 0.06179 (0.06821) 0.1708 (0.0657) 0.2991 (0.3289) 1.042 (1.746)
0.6 Time 0.5028 (0.1514) 0.1376 (0.1614) 0.05767 (0.02264) 0.05976 (0.02864) 0.1421 (0.04185) 0.1573 (0.2019) 0.3919 (0.93)
0.7 Time 0.3975 (0.1663) 0.1015 (0.1064) 0.07437 (0.07595) 0.08195 (0.1162) 0.1253 (0.03575) 0.1538 (0.2255) 0.3738 (1.007)
0.8 Time 0.3499 (0.1075) 0.09129 (0.03348) 0.07957 (0.03062) 0.07738 (0.03206) 0.142 (0.03733) 0.1259 (0.1176) 0.1909 (0.06753)
0.9 Time 0.3183 (0.1172) 0.0907 (0.03449) 0.09587 (0.1122) 0.1031 (0.1207) 0.1454 (0.02508) 0.1609 (0.2382) 0.2925 (0.9214)
1.0 Time 0.303 (0.05565) 0.1067 (0.0494) 0.1029 (0.0561) 0.1073 (0.0597) 0.1552 (0.03678) 0.1841 (0.2233) 0.1944 (0.06617)
Table 10: Mean Time (SD) for $d=128,s=10$ and $\sigma=0.05$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 Time 0.01993 (0.01462) 0.07793 (0.06112) 0.00739 (0.00417) 0.008899 (0.005657) 0.08419 (0.06419) 0.1891 (0.09824) 0.2951 (0.1907)
0.2 Time 0.1336 (0.04051) 0.3351 (0.1643) 0.02581 (0.01689) 0.02901 (0.0108) 0.4944 (0.2485) 0.61 (0.1112) 1.245 (0.4835)
0.3 Time 0.3753 (0.1275) 0.2123 (0.1632) 0.0257 (0.01092) 0.04378 (0.08122) 0.302 (0.2415) 0.4586 (0.2208) 1.296 (1.137)
0.4 Time 0.6192 (0.1141) 0.1942 (0.1632) 0.03955 (0.02311) 0.04445 (0.0207) 0.191 (0.1407) 0.4549 (0.3075) 1.53 (1.698)
0.5 Time 0.6784 (0.1806) 0.1867 (0.179) 0.05155 (0.02586) 0.06454 (0.1035) 0.1413 (0.05757) 0.3199 (0.3196) 0.9733 (1.625)
0.6 Time 0.4605 (0.161) 0.1458 (0.1995) 0.04942 (0.02432) 0.05805 (0.06407) 0.1279 (0.04818) 0.2552 (0.2949) 1.105 (2.237)
0.7 Time 0.3264 (0.1112) 0.0944 (0.1296) 0.05519 (0.02142) 0.06366 (0.08287) 0.1171 (0.04002) 0.1778 (0.1704) 0.33 (0.7549)
0.8 Time 0.3621 (0.118) 0.1184 (0.1856) 0.0924 (0.08049) 0.09844 (0.07525) 0.1592 (0.04999) 0.1901 (0.1602) 0.3504 (1.017)
0.9 Time 0.3463 (0.1191) 0.1029 (0.05027) 0.1115 (0.1813) 0.1292 (0.215) 0.1702 (0.03466) 0.2192 (0.1908) 0.2837 (0.6244)
1.0 Time 0.3231 (0.1258) 0.09973 (0.035) 0.1363 (0.2276) 0.1741 (0.3248) 0.1673 (0.04315) 0.2882 (0.3655) 0.2437 (0.1679)
Table 11: Mean Time (SD) for $d=256,s=15$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 Time 0.05887 (0.01293) 0.2234 (0.08841) 0.01697 (0.005035) 0.01699 (0.004914) 0.1988 (0.09614) 0.3827 (0.07207) 0.6776 (0.1144)
0.2 Time 0.1291 (0.04144) 0.3178 (0.1516) 0.01979 (0.004716) 0.02326 (0.01892) 0.2496 (0.1418) 0.4005 (0.08792) 2.426 (1.495)
0.3 Time 0.7639 (0.2156) 0.289 (0.2028) 0.04582 (0.01895) 0.05365 (0.02381) 0.2879 (0.1784) 0.7338 (0.3489) 4.571 (3.882)
0.4 Time 1.514 (0.2979) 0.4177 (0.3474) 0.1032 (0.06255) 0.123 (0.1618) 0.3198 (0.1482) 0.7819 (0.6811) 4.638 (5.593)
0.5 Time 1.508 (0.4161) 0.3399 (0.3124) 0.1315 (0.05244) 0.1227 (0.03739) 0.291 (0.09094) 0.5467 (0.6743) 2.529 (4.824)
0.6 Time 3.892 (1.556) 1.135 (1.459) 0.5243 (0.1609) 0.5964 (0.9066) 1.018 (0.2602) 1.366 (1.971) 3.584 (9.087)
0.7 Time 2.256 (1.078) 0.6103 (0.5344) 0.4549 (0.08223) 0.4341 (0.0722) 0.7783 (0.1122) 0.9275 (1.521) 2.688 (9.078)
0.8 Time 1.738 (0.2359) 0.5042 (0.092) 0.5566 (0.7428) 0.5519 (0.7202) 0.7686 (0.1213) 0.702 (0.708) 1.018 (0.9496)
0.9 Time 1.576 (0.1965) 0.531 (0.1265) 0.7153 (0.9071) 0.73 (1.01) 0.7896 (0.1359) 1.044 (1.544) 0.9093 (0.4753)
1.0 Time 1.438 (0.183) 0.5723 (0.2167) 0.645 (0.6552) 0.6464 (0.5954) 0.7828 (0.09955) 0.8698 (0.9806) 0.9996 (0.9732)
Table 12: Mean Time (SD) for $d=256,s=15$ and $\sigma=0.05$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 Time 0.06015 (0.01408) 0.2368 (0.08766) 0.01761 (0.004725) 0.01972 (0.01006) 0.2015 (0.08702) 0.3895 (0.06751) 0.6818 (0.07771)
0.2 Time 0.1358 (0.06205) 0.3425 (0.1847) 0.02148 (0.008248) 0.02143 (0.007374) 0.2777 (0.1744) 0.4298 (0.1278) 2.522 (1.414)
0.3 Time 0.859 (0.2789) 0.3643 (0.2975) 0.04648 (0.02163) 0.06987 (0.1505) 0.2603 (0.182) 0.7928 (0.3222) 5.155 (3.626)
0.4 Time 1.832 (0.2002) 0.484 (0.3226) 0.1184 (0.06135) 0.1091 (0.04349) 0.3176 (0.09057) 0.8818 (0.7331) 4.047 (5.325)
0.5 Time 1.812 (0.435) 0.3944 (0.3735) 0.1328 (0.04252) 0.1257 (0.03247) 0.2781 (0.08556) 0.6853 (0.7801) 2.898 (5.014)
0.6 Time 1.234 (0.4237) 0.3135 (0.3604) 0.1599 (0.04195) 0.1796 (0.2797) 0.2856 (0.06056) 0.4151 (0.5578) 1.49 (4.121)
0.7 Time 0.9523 (0.2734) 0.2349 (0.275) 0.1719 (0.0329) 0.1647 (0.03235) 0.2858 (0.06317) 0.3363 (0.3792) 0.654 (2.279)
0.8 Time 0.7881 (0.1328) 0.2051 (0.05422) 0.1973 (0.1323) 0.201 (0.1327) 0.2805 (0.04738) 0.3521 (0.2724) 0.4649 (0.4407)
0.9 Time 0.7322 (0.1035) 0.2063 (0.02253) 0.2012 (0.03551) 0.2529 (0.3343) 0.2927 (0.04338) 0.3249 (0.1853) 0.4405 (0.2648)
1.0 Time 0.6781 (0.0835) 0.2296 (0.05449) 0.2365 (0.1932) 0.2358 (0.1469) 0.3107 (0.05463) 0.4408 (0.4755) 0.4186 (0.06463)